### v3

# FPT Parameterisations of Fractional and Generalised Hypertree Width

Abstract.
We present the first fixed-parameter tractable (FPT) algorithms for exact computation of generalized hypertree width ($\mathsf{ghw}$) and fractional hypertree width ($\mathsf{fhw}$). Our algorithms are parameterized by the target width, the rank, and the maximum degree of the input hypergraph. More generally, we show that testing $f$-width is in FPT for a broad class of width functions that we call manageable. This class contains the edge cover number $\rho$ and its fractional relaxation $\rho^{*}$, and thus covers both generalized and fractional hypertree width. We additionally extend our framework to also obtain an fpt algorithm for computing a discretized version of adaptive width. Our approach extends a recent algorithm for treewidth (Bojańcyk & Pilipczuk, LMCS 2022) that utilises monadic second-order transductions.
To extend this idea beyond treewidth we develop new combinatorial machinery around elimination forests in hypergraphs, culminating in a structural normal form for optimal witnesses that makes transduction-based optimisation applicable in the much more general context of manageable width functions.
This yields the first exact FPT algorithms for these measures under any nontrivial parameterisation and provides structural tools that may enable more direct optimisation algorithms.
hypertree decompositions, generalised hypertree width, fractional hypertree width, fixed parameter tractable
copyright: ccjournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 2 (PODS)article: 104publicationmonth: 5doi: 10.1145/3801900isbn: 978-1-4503-XXXX-X/2018/06ccs: Theory of computation Fixed parameter tractabilityccs: Theory of computation Database query processing and optimization (theory)
1. Introduction
A tree decomposition of a hypergraph $H$ is a tree $T$ with a function $\mathsf{bag}$ which assigns each vertex in $T$ a set of vertices of $H$ such that certain conditions hold. The $f\textsf{-width}$ of a tree decomposition $(T,\mathsf{bag})$ is defined as $\max_{u\in V(T)}f(\mathsf{bag}(u))$. This generalises various widely studied (hyper)graph parameters in the literature. For instance, when $f$ is the function $B\mapsto|B|-1$, the $f\textsf{-width}$ is treewidth (Robertson and Seymour, 1986), one of the most prominent concepts in the study of graph algorithms. When $f$ is the edge cover number $\rho$ of the bag, or its fractional relaxation $\rho^{*}$, the $f\textsf{-width}$ equals generalized hypertree width ($\mathsf{ghw}$(Gottlob et al., 2007) and fractional hypertree width ($\mathsf{fhw}$(Grohe and Marx, 2014), respectively. These parameters have been shown central to the complexity of a wide range of problems whose instances are naturally represented as hypergraphs, e.g., conjunctive query evaluation (Gottlob et al., 2002; Atserias et al., 2013; Ngo et al., 2018), graph and hypergraph counting problems (Gottlob et al., 2021; Bressan et al., 2023, 2025), problems in game theory (Gottlob et al., 2005a), or combinatorial auctions (Gottlob and Greco, 2013).
[…]
Recently significant progress has been made on identifying computationally easier fragments of these problems. Gottlob et al. (Gottlob et al., 2021, 2020) showed that deciding whether $\mathsf{ghw}(H)\leq k$ and $\mathsf{fhw}(H)\leq k$ is in the complexity class XP when parameterised by the width $k$, i.e., permitting an $O(n^{f(k)})$ algorithm, under loose structural restrictions on $H$. Even more recently, Lanzinger and Razgon (Lanzinger and Razgon, 2024) gave the first FPT approximation algorithm for $\mathsf{ghw}$ for hypergraphs with bounded edge-intersection cardinality and in parallel Korchemna et al. (Korchemna et al., 2024) showed that even polynomial time approximation of $\mathsf{ghw}$ and $\mathsf{fhw}$ is possible under similar structural restrictions.
Nonetheless, as of now, no fixed-parameter tractable algorithms under any non-trivial parametrisations are known for either problem.
In this paper we present the first exact FPT algorithms for deciding $\mathsf{ghw}(H)\leq k$ and $\mathsf{fhw}(H)\leq k$. Since the natural parametrization by width $k$ alone is known to be paraNP-hard, obtaining FPT algorithms requires parameterising beyond width. Concretely, we additionally consider the rank (the maximum edge cardinality) and the maximum vertex degree (denoted by $\Delta(\cdot)$) of the input hypergraph as parameters. Formally we study the following class of problems.
       $f$-width-check
Input Hypergraph $H$ and $k>0$.
Parameters $\mathit{rank}(H)$, $\Delta(H)$ and $k$
Output $f\textsf{-width}(H)\leq k$?
In order to solve this question for $\mathsf{ghw}$ and $\mathsf{fhw}$, we show that deciding $f\textsf{-width}$ in general is FPT in this setting for a large class of functions $f$ that we will call manageable. Informally, a manageable $f$ is additive on non-adjacent disjoint sets and $f(U)\leq k$ implies bounds on the size of $U$, in addition to some minor technical conditions (the notion is introduced formally in Section˜3).
In particular, both $\rho$ and $\rho^{*}$ are manageable width functions and thus $\mathsf{ghw}$ and $\mathsf{fhw}$ are instances of $f\textsf{-width}$ with manageable $f$. Our main result then is the following.
Theorem 1.
$f$-width-check is fixed parameter tractable for any manageable width function $f$.
We approach this problem by building on the transduction-based treewidth algorithm of Bojańczyk and Pilipczuk (Bojańczyk and Pilipczuk, 2022). At a high level, their result shows how to express the process of turning a potentially suboptimal tree decomposition into an optimal one by an MSO transduction, based on deep combinatorial insights on the fine structure of so-called elimination forests (see Section˜2). In our setting, we follow the same blueprint: we construct an MSO transduction that, given an appropriate bounded-treewidth backbone, produces a candidate decomposition and an MSO test that the produced decomposition has $f$-width at most $k$.
[…]
Our main technical contribution is the extension of the approach of (Bojańczyk and Pilipczuk, 2022) to width functions that only permit such weaker compositionality for edge-disjoint partitions. This requires large-scale changes from the treewidth result, combining multiple additional combinatorial insights with explicit abstraction of various implicit structural observations. A secondary motivation for isolating this structure is that it is not tied to MSO. Our combinatorial analysis concludes in a hypergraph dealternation principle (Reduced Hypergraph Dealternation) that shows that optimal witnesses are in a certain technical sense structurally well-behaved. As argued for treewidth in (Bojańczyk and Pilipczuk, 2022), such principles are exactly what underlies Bodlaender-Kloks-style direct FPT algorithms for treewidth. We therefore view our results as providing not only an MSO/transduction-based FPT route, but also a concrete combinatorial foundation for more direct and practical FPT algorithms for $\mathsf{ghw}/\mathsf{fhw}$.
As a final result, our framework produces an MSO test for $f$-width $\leq k$. This lets us go beyond a single $f$ and handle certain families of width functions in one shot. As an illustration, we consider a discretized variant of adaptive width (Marx, 2011), defined as the $\mathcal{F}$-width where $\mathcal{F}_{\delta}$ ranges over independent-set functions whose vertex weights are multiples of $1/\delta$ (and every hyperedge has total weight at most $1$). We show that this discretized adaptive width can be computed exactly in FPT time.
Theorem 2.
For each integer $\delta\geq rank(H)$, $\mathcal{F}_{\delta}$-width is fixed-parameter tractable parameterised by $k$, $\mathit{rank}(H)$, $\Delta(H)$, and $\delta$.
Theorem 2 leaves an interesting open question as to whether the discretized adaptive width can approximate the actual adaptive width. We provide some initial insight related to this question. We consider a relaxed version $\mathcal{F}^{*}_{\delta}$ of $\mathcal{F}_{\delta}$ consisting of all independent set functions where all the non-zero weights are at least $\delta$ (or, to put it differently, there are no small non-zero weights $0<x<\delta$). We show that Theorem 2, combined with a simple rounding strategy, implies factor $2$ FPT approximation of $\mathcal{F}^{*}_{\delta}$-width.
The rest of this paper is structured as follows. We formally define tree decompositions and key width measures, as well as further necessary technical preliminaries in Section˜2. We present an extended technical overview of our main results and the central statements leading to it in Section˜3. The lifting of our framework to sets of functions and the corresponding results for adaptive width are presented in Section˜4. Closing remarks and discussion of a number of interesting new combinatorial open problems raised by our results are given in Section˜5. To improve presentation some further proof details follow in the appendix.
2. Preliminaries
We say a family $(U_{i})_{i\in I}$ of sets is a weak partition of a set $V$ if $\cup_{i\in I}U_{i}=V$ and $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$. If no $U_{i}$ is empty, we say it is a partition of $V$.
Graphs & Hypergraphs    A hypergraph $H$ is a tuple $(V,E)$ where $V$ is a finite set and $E$ is a set of non-empty subsets $e\subseteq V$. If all $e\in E$ have exactly two elements, we refer to $H$ as a (simple) graph. We call $V$ the vertices of $G$ and $E$ the edges of $H$. For a hypergraph $H$, we denote the vertices of $H$ by $V(H)$ and the edges by $E(H)$. We write $I(v):=\{e\in E(H)\mid v\in e\}$ for the set of edges incident to vertex $v$, and we say two vertices $u\neq v$ are connected if $I(u)\cap I(v)\neq\emptyset$. For a vertex subset $U\subseteq V(H)$, we call the hypergraph $H^{\prime}$ with $V(H^{\prime})=U$ and $E(H^{\prime})=\{U\cap e:e\in E(H),~U\cap e\neq\emptyset\}$ the subgraph of $H$ induced by $U$, denoted by $H[U]$. For graphs, we only keep edges of cardinality $2$ in the induced subgraph, i.e. only edges contained in $U$. We assume throughout that hypergraphs have no isolated vertices, i.e., every vertex in a hypergraph is contained in some edge.
Let $G$ be a graph. A path from $v_{1}$ to $v_{\ell}$ in $G$ is a sequence $(v_{1},v_{2},\cdots,v_{\ell})$ of distinct vertices such that $\{v_{i},v_{i+1}\}\in E$ for $1\leq i\leq\ell-1$. A cycle is a path except that $v_{1}=v_{\ell}$. A graph is connected if, for any $v_{i},v_{j}\in V$, there is a path from $v_{i}$ to $v_{j}$. A connected graph without cycles is a tree, and a disjoint union of trees is called a forest.
A rooted tree is a tree $T$ together with a designated vertex $r\in V(T)$ called the root. For a vertex $v\in V(T)$, we refer to the vertices on the unique path from $r$ to $v$ as the ancestors of $v$. If $v_{1}$ is an ancestor of $v_{2}$, we refer to $v_{2}$ as a descendant of $v_{1}$. We will denote the subtree of $T$ induced by the set of all descendants of a vertex $x$ in $T$ as $T_{x}$. A disjoint union of rooted trees is a rooted forest $F$, and we refer to the roots of the trees as the roots of $F$. Note that this implies that every vertex is both ancestor and descendant to itself.
Hypergraph Decompositiosn
A tree decomposition of hypergraph $H$ is a tree $T$ together with a function $\mathsf{bag}:T\to 2^{V(H)}$ such that, for every edge $e\in E(H)$, there is a vertex $n\in V(T)$ such that $e\subseteq\mathsf{bag}(n)$
. Additionally, we require that the subgraph (of $T$) induced by $\{n\in V(T):v\in\mathsf{bag}(n)\}$ is connected for each $v\in V(H)$. We will refer to the set $\{\mathsf{bag}(n):n\in T\}$ as the bags of $T$. Throughout the paper we will consider tree decompositions to be rooted. Note that this is not a technical restriction as any arbitrary rooting is sufficient.
A width function $f$ is a function that assigns a hypergraph $H$ together with a subset $U\subseteq V(H)$ of its vertices a non-negative real number $f(H,U)$. We will commonly write this as $f_{H}(U)$, and, if the underlying hypergraph is clear from context, simply as $f(U)$. For a given tree decomposition $(T,\mathsf{bag})$ of $H$, the $f\textsf{-width}$ of $(T,\mathsf{bag})$ is the maximal width of any bag, i.e. $\max_{v\in T}f(H,\mathsf{bag}(v))$. The $f\textsf{-width}$ of $H$ is the minimal $f\textsf{-width}$ over all tree decompositions of $H$. We denote this as $f\textsf{-width}(T)$ and $f\textsf{-width}(H)$ respectively. Now let $H$ be a hypergraph and $U\subseteq V(H)$. We say a function $\lambda:E(H)\to\{0,1\}$ is an edge cover of $U$ if $\sum_{\begin{subarray}{c}e:v\in e\end{subarray}}\lambda(e)\geq 1$ for all $v\in U$. In other words, we choose a set of edges such that each vertex is included in at least one of them. The weight of an edge cover is $\sum_{e\in E}\lambda(e)$, and the edge cover number $\rho(H,U)$ is the minimal weight of an edge cover of $U$. The generalized hypertreewidth ($\mathsf{ghw}$) of a hypergraph $H$ is the $f\textsf{-width}(H)$ where $f=\rho$. The fractional cover number $\rho^{*}(H,U)$ is defined analogously, except that we can assign non-integer values to edges, i.e. $\lambda:E(H)\to[0,1]$. The fractional hypertreewidth ($\mathsf{fhw}$) of a hypergraph $H$ is the $f\textsf{-width}(H)$ where $f=\rho^{*}$.
Factors and Elimination Forests    Our technical development involves extensive combinatorial arguments about the structure of certain rooted forests. In particular we argue on the structure in terms of over so-called factors as introduced in (Bojańczyk and Pilipczuk, 2022).
Definition 1 (Factors).
For rooted forest $T$ we define the following three kinds of factors:
Tree Factor:
A tree factor is a set $V(T_{x})$ for some $x\in V(T)$. We call $x$ the root of the factor.
Forest Factor:
A forest factor is a nonempty union of tree factors whose roots are siblings. In particular, a tree factor is also a forest factor but the empty set is not. The set of roots of the tree factors included in the union are called the roots of the
factor. If they are not roots of components of $T$, we define their joint parent as the parent of the factor. Note that We consider two nodes of $T$ siblings when they have a joint parent or if they both roots of components of $T$. This convention is chosen to remain consistent with the notation of (Bojańczyk and Pilipczuk, 2022).
Context Factor:
A context factor is a set $U=X\setminus Y$ where $X$ is a tree factor and $Y$ is a forest factor whose roots are strict descendants of the root of $X$. The root of $U$, denoted by $rt(U)$, is the root of $X$ and the roots of $Y$ are called the appendices of $U$. Note that a tree factor is not a context factor. We will sometimes refer to $X$ as the base tree factor of $U$ and to $Y$ as the removed forest of $U$. For clarity we sometimes denote $X$ and $Y$ as $X(U)$ and $Y(U)$. We also denote the parent of the appendices by $pr(U)$ and let the spine of $U$, denoted by $spine(U)$, be the path from $rt(U)$ to $pr(U)$
We say that two factors are intersecting if there have at least one vertex in common. Note also that a tree factor is also always forest factor.
Let $X$ be a set of nodes of a forest $T$, an $X$-factor (wrt. $T$) is a factor $F$ with $F\subseteq X$. A factorization of $X$ is a partition of $X$ into $X$-factors. A $X$-factor is maximal if no other $X$-factor contains it as a strict subset. Let $\mathit{MF}(X)$ denote the set of maximal $X$-factors. Note that by Lemma 3.2 (Bojańczyk and Pilipczuk, 2022) we have that $\mathit{MF}(X)$ is always a factorization of $X$.
We will be interested in particular in factors of elimination forests. Intuitively, this is a tree-shaped arrangement of the vertices of a hypergraph that takes adjacency into account. We will see that this representation also naturally induces tree decompositions.
Definition 2 (Elimination Forest).
Let $G$ be a hypergraph. An elimination forest $T$ for $G$ is a rooted tree over $V(G)$ such that $u,v\in e$ for some $e\in E(G)$ implies that either $u$ is an ancestor of $v$ in $T$ or $v$ is an ancestor of $u$.
An elimination forest $T$ induces a function $\mathsf{bag}_{T}:$ $V(T)\to 2^{V(G)}$ as follows: $\mathsf{bag}_{T}(u)$ contains exactly $u$ and all those ancestors $v$ that are adjacent with $u$ or a descendant of $u$. Note that $(T,\mathsf{bag}_{T})$ is a tree decomposition of $H$. In what follows, when we refer to the $f\textsf{-width}$ of an elimination forest, we mean the $f\textsf{-width}$ of the tree decomposition induced by the forest.
$v_{1}$$v_{3}$$v_{4}$$v_{5}$$v_{2}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$\{v_{1}\}$$\{v_{1},v_{2}\}$$\{v_{1},v_{3}\}$$\{v_{1},v_{3},v_{4}\}$$\{v_{3},v_{5}\}$
Figure 1. An elimination forest for the depicted hypergraph (left) and its induced tree decomposition (right).
Tree decompositions induced by elimination forests will play a central role in our technical developments. We will now fix some auxiliary definitions to simplify reasoning over tree decompositions. Note that while tree decompositions in this paper will generally be induced by an elimination forest, the definitions are independent of the forest and apply generally. To that end, for tree decomposition $\mathbf{t}=(T,B)$ we define the following.
For a non-root node $x$ of $T$ with parent $p$, the adhesion is $adh(x):=\mathsf{bag}(x)\cap\mathsf{bag}(p)$. The adhesion of the root node is the empty set.
The margin of a node $x$ of $T$ are those vertices of $\mathsf{bag}(x)$ that are not in the adhesion of $x$, i.e., $mrg(x):=\mathsf{bag}(x)\setminus adh(x)$.
Finally, the component of node $x$ of $T$ is $cmp(x):=\bigcup_{y\in V(T_{x})}mrg(y)$.
The target sets of $\mathbf{t}$ are the components of nodes of $T$. That is $\mathsf{TS}(\textbf{t})=\{cmp(x)\mid x\in V(T)\}$.
When there is an underlying elimination forest that is not clear from the description, we state it in the subscript: say $adh_{T}(x)$.
xxadh(x):adh(x):mrg(x):mrg(x):$cmp(x):$
(a) The adhesion, margin and component of the node $x$.
(b) A Tree, Forest and Context Factor. Roots marked in white, appendices in orange.
Figure 2. Illustrations of factors and key tree decomposition notions used in this paper.
MSO Transductions    Ultimately our FPT algorithm will be stated in terms of MSO transductions.
A transduction is a mapping of a model over the input signature to a set of models over the output signature that is defined as a composition of elementary transductions of Colouring, Copying, MSO Interpretation, Filtering, and Universe restriction. (a detailed description of elementary transductions is provided Appendix D). In particular, let ${\bf g}_{1},\dots\bf g_{p}$ be a sequence of elementary transductions such that for each $2\leq i\leq p$ the input signature of ${\bf g}_{i}$ is the same as the output signature of ${\bf g}_{i-1}$.
Then the transduction $Tr[{\bf g}_{1},\dots,{\bf g}_{p}]$ is a function over the models $M$ of the input signature of ${\bf g}_{1}$ defined as follows. If $p=1$ then $Tr[{\bf g}_{1}](M)={\bf g}_{1}(M)$. Otherwise $Tr[{\bf g}_{1},\dots,{\bf g}_{p}](M)=\bigcup_{M^{\prime}\in{\bf g}_{1}(M)}Tr[{\bf g}_{2},\dots,{\bf g}_{p}](M^{\prime})$. The size of an MSO transduction is the sum of the sizes of its elementary MSO transductions, where colouring has size 1, copying has as size the number $k$ of copies it produces, and for any other type of elementary transduction the size is the total size of MSO formulas (i.e., number of symbols) that make up its description. We write $\|\mathbf{Tr}\|$ for the size of transduction $\mathbf{Tr}$.
We note that in (Bojańczyk and Pilipczuk, 2022), a transduction is defined as a binary relation with pairs $(M_{L},M_{R})$ where the ’left-hand’ model $M_{L}$ is over the input signature whereas the ’right-hand’ model $M_{R}$ is over the output signature. Our definition is equivalent as a mapping of a left-hand model to a set of right model determines a binary relation and vice versa. For a detailed description of MSO logic and formal languages we refer the reader to (Courcelle and Engelfriet, 2012).
We will state MSO formulas over two specific signatures, $\tau_{td}$ for formulas over tree decompositions, and $\tau_{ef}$ for formulas over elimination forests. Formally, the signature $\tau_{td}$ consists of unary predicates $\mathit{Vertex},\mathit{Edge},\mathit{Node}$ and binary predicates $\mathit{Bag},\mathit{Adjacenct},\mathit{Descendant}$. The signature $\tau_{ef}$ consists of unary predicates $\mathit{Vertex},\mathit{Edge}$ and binary relations $\mathit{Child},\mathit{Adjacenct}$.
The intended meaning of these predicates is the following: for a model $M$ over $\tau_{td}$, $\mathit{Vertex}$ and $\mathit{Edge}$ are the vertices and (hyper)edges of a hypergraph $H(M)$, and $\mathit{Adjacent}$ describes, for a given tuple $(v,e)$, if vertex $v$ is contained in the edge $e$. NodesNodes and $Descendant$ describe the vertices and descendancy of a tree $T$ and the predicate $Bag$ induces a bag function bagbag (via $bag(t)=\{v\in\mathit{Vertex}\mid Bag(t,v)\}$) such that ${\bf t}=(T,bag)$ is a tree decomposition of $H(M)$. For a hypergraph $H$ and a tree decomposition ${\bf t}=(T,{\bf B})$, we define the model $M=M(H,{\bf t})$ induced by $H$ and $\bf t$} as the model $M$ with $H=H(M)$ and $\bf t$} being the tree decomposition described by $Node$, $\mathit{Descendant}$ and $Bag$ relation of $M$.
For a model $M$ over $\tau_{ef}$, $\mathit{Vertex},Edge$ and $\mathit{Adjacent}$ describe a hypergraph $H(M)$ (as with $\tau_{td}$), while the predicates $\mathit{Vertex}$ and $\mathit{Child}$ model a directed graph $F(M)$ that is, in fact, will be proved to be an elimination forest of $H(M)$ if $M$ belongs to the output of the transduction being constructed.
3. Technical Overview
To show our main result, we require a substantial amount of rather technical combinatorial arguments. In this section we therefore focus on providing a technical overview, presenting the main combinatorial results, how they are derived, and how they interact to form the final FPT algorithm. We first define the notion of a manageable width function in Section˜3.1 In Section˜3.2 and Section˜3.3, we develop the main technical ingredients for the proof of the theorem and provide the proof in Section˜3.4. Full proof details are provided in Appendix sections A to D.
3.1. The Main Theorem
We isolate the minimal properties of a width function that we need in order to extend the transduction-based optimisation framework of Bojańczyk and Pilipczuk (Bojańczyk and Pilipczuk, 2022). We call such width functions manageable. This class includes the edge-cover number $\rho$ and its fractional relaxation $\rho^{\ast}$, and thus subsumes generalized and fractional hypertree width. At the same time, it is broad enough to plausibly capture further hypergraph width measures studied in the literature (e.g., variants based on modular width functions (Marx, 2011)).
A key obstacle compared to treewidth is that the relevant width functions are not compositional with respect to arbitrary partitions of a bag. We therefore distinguish a strong and a weak notion of separability.
Definition 1 (Monotone separability).
[…]
$f_{H}(U)=g_{f}\bigl(f_{H}(U_{1}),f_{H}(U_{2})\bigr).$
[…]
The treewidth algorithm of (Bojańczyk and Pilipczuk, 2022) crucially relies on the underlying width function being monotone separable. Our main technical step is to replace it by weak separability, which requires substantial technical extension and additional structure to control how decompositions interact with edge-disjoint splits. We capture the extra requirements in the following definition.
Definition 2.
We call a width function $f$ manageable if it satisfies all of the following.
Weak monotone separability.:
$f$ is weakly monotone separable, as in Definition 1.
Bounded size.:
There exists a computable function $\beta$ such that for every $k\geq 0$, every hypergraph $H$, and every $U\subseteq V(H)$, if $f_{H}(U)\leq k$, then $|U|\leq\beta\bigl(k,\mathit{rank}(H)\bigr).$ We refer to any such $\beta$ as a size bound for $f$.
Invariant locality.:
For all hypergraphs $H_{1},H_{2}$ and all $U_{1}\subseteq V(H_{1})$, $U_{2}\subseteq V(H_{2})$, if $H_{1}[U_{1}]\cong H_{2}[U_{2}]$, then $f_{H_{1}}(U_{1})=f_{H_{2}}(U_{2})$.
Automatizability.:
There exists an algorithm that, given a hypergraph $H$, computes $f_{H}(V(H))$ in time at most $t(|V(H)|)$ for some computable function $t$.
Intuitively, invariant locality means that $f_{H}(U)$ depends only on the isomorphism type of the induced subhypergraph $H[U]$.
Together with bounded size, this yields a bounded number of values phenomenon: for fixed $k$ and $r$, there are only finitely many isomorphism types of induced subhypergraphs $H[U]$ with $\mathit{rank}(H)\leq r$ and $f_{H}(U)\leq k$, and hence only finitely many values $\leq k$ that $f$ can take on such bags (bounded by a function of $k$ and $r$). Finally, automatizability is necessary only at the final stage, to effectively enumerate (up to isomorphism) all small hypergraphs of f\!f-width at most $k$.
We are now in a position to state our main result regarding exact computation of $f\textsf{-width}$ in full formality.
Theorem 3 (Extended version of Theorem˜1).
$f$-width-check is fixed-parameter tractable for every manageable width function $f$. In particular, there is an $g(k,\mathit{rank}(H),\Delta(H)\cdot|V(H)|)$, for some computable function $g$, time algorithm that takes as input a hypergraph $H$ and a parameter $k$ and returns a tree decomposition $H$ with width $f\textsf{-width}(H)$ if $f\textsf{-width}(H)\leq k$, or rejects. In the latter case it is guaranteed that $f\textsf{-width}(H)>k$.
Recall, $\mathsf{ghw}$ is precisely $f\textsf{-width}$ when $f$ is the edge cover number $\rho$, and $\mathsf{fhw}$ is $f\textsf{-width}$ when $f$ is the fractional edge cover number $\rho^{*}$. It is trivial to verify that $\rho$ is indeed manageable. For $\rho^{*}$, the bounded size follows from
(1) $|U|\leq\sum_{v\in U}\underbrace{\sum_{e:v\in e}\lambda(e)}_{\geq 1}=\sum_{e\in E}\sum_{v\in e}\lambda(e)\leq\sum_{e\in E}\mathit{rank}(H)\lambda(e)\leq k\cdot\mathit{rank}(H).$
The other conditions of manageability are immediate.
The long-standing open question of fixed-parameter tractable algorithms for $\mathsf{ghw}$ and $\mathsf{fhw}$ then directly follow from Theorem˜3.
Corollary 4.
For input hypergraph $H$ and $k>0$, it is fixed-parameter tractable to check $\mathsf{ghw}(H)\leq k$ and $\mathsf{fhw}(H)\leq k$ parameterised by $k,\mathit{rank}(H),\Delta(H)$.
In the rest of this section, we will provide an overview of the proof of Theorem˜3. Subsequent sections will then expand on the technical details of the statements made here.
3.2. The Structure of Optimal Elimination Forests
In this subsection, we develop the main combinatorial engine behind Theorem 3, which we call the Reduced Hypergraph Dealternation Lemma. The key to our algorithm is that we can check manageable $f\textsf{-width}$ on elimination forests whose structure is simple. However, this simplicity is relative in the following sense. Let ${\bf t}=(T,{\bf B})$ be a tree decomposition of $H$ whose $f$-width is not much larger than $f\textsf{-width}(H)$. Then there exists an optimal elimination forest $F$ for $H$ (i.e., $f\textsf{-width}(F)=f\textsf{-width}(H)$) whose induced decomposition is aligned with $\bf t$}.
This alignment in particular refers to how factorisations of the elimination forest match up with $\bf t$}. We will therefore be interested in measures that relate the factorisation of elimination forests to specific tree decompositions. In the following, we refer to a pair of a tree decomposition and an elimination forest for the same hypergraph $H$ as a $\mathit{TF}$-pair.
We introduce two new measures to quantify the relevant structural alignment of $\mathit{TF}$-pairs. The first factor in such efficient computability is that we need each target set of $\bf t$} to have a small factorization in $F$. Intuitively, this restricts the possible ways to construct $F$ from $\bf t$} (recall, target sets $\mathsf{TS}$ and maximal factors $\mathit{MF}$ are defined in Section˜2) .
Definition 5.
For $\mathit{TF}$-pair ${\bf t},F$, we define the split of the pair as the size of the largest maximal factorization of target sets of $\bf t$, i.e., $\mathit{split}({\bf t},F):=\max_{U\in\mathsf{TS}({\bf t})}|\mathit{MF}(U)|$.
The second aspect to this structural alignment is that the individual factors themselves follow a kind of top-down monotonicity. Formally, this means that we want to avoid the presence of many context-factors in our representation. For the definition, it will be helpful to extend our notation. We call $U\subseteq V(H)$ well-formed w.r.t. $F$ if $\mathit{MF}(U)$ contains no context factor (equivalently, it consists only of forest factors).
Definition 6.
For $\mathit{TF}$-pair ${\bf t},F$ we define the irregularity of a node $x$ of $\bf t$} as the number of children $y$ of $x$, such that $cmp(y)$ is not well-formed w.r.t. $F$. We denote the maximum irregularity of a node of $T$ by $\mathit{mir}({\bf t},F)$.
These two measures of structural complexity, $\mathit{split}$ and $\mathit{mir}$, are key to our framework. In key contrast to (Bojańczyk and Pilipczuk, 2022), we isolate split and irregularity as the only interface parameters necessary to build a modular, general theoretical framework for width checking using MSO transductions. It allows us to decouple various statements for treewidth into independent parts, connected only through split and irregularity. This abstraction and generalisation ultimately allows us to establish core properties also for width functions that are not monotone separable.
Lemma 7 (Reduced Hypergraph Dealternation).
Assume that $f$ is manageable. There exists a monotone function $\gamma$ such that for every hypergraph $H$ and every tree decomposition $\bf t$} of $H$, there is an elimination forest $F$ of $H$ with $f\textsf{-width}(F)=f\textsf{-width}(H)$ and
$\mathit{split}({\bf t},F),\ \mathit{mir}({\bf t},F)\ \leq\ \gamma\bigl(\mathsf{tw}({\bf t}),\,f\textsf{-width}(H),\,\mathit{rank}(H),\,\Delta(H)\bigr).$
Proof Sketch.
We analyse partitions $(Red,X,Blue)$ of $V(H)$, where $X$ separates $Red$ from $Blue$ (i.e., no hyperedge has both a red and a blue vertex). In the application, $X$ is chosen as a subset of a bag $B_{x}$ of $\bf t$}, and hence $|X|\leq\mathsf{tw}({\bf t)}+1$. We further consider only reduced elimination forests: an elimination forest $F$ is reduced if every parent–child edge $(u,v)$ of $F$ is witnessed by some hyperedge, i.e., there exists $e\in E(H)$ with $u\in e$ and $e\cap V(F_{v})\neq\emptyset$, where $F_{v}$ is the subtree rooted at $v$. We show that one can transform every elimination forest into a reduced one of the same width, hence this restriction can be made without loss of generality.
Fix a reduced optimal elimination forest $F$ and a coloured context factor $U\subseteq Red\cup Blue$. Along the spine of $U$, consider the maximal monochromatic subpaths (intervals). The number of such intervals is a coarse measure of how often the spine alternates between the two sides of the cut, and thus of how “entangled” this factor is with the separator $X$: many intervals indicate a complicated interaction that we will later rule out in a well-aligned optimal forest. A swap exchanges two consecutive intervals along the spine. In the alternating pattern of four consecutive intervals, swapping the middle two makes adjacent intervals merge, reducing the total number of intervals by 2 while preserving connectivity and monochromaticity (cf. Figure˜3).
$P_{i-1}$$P_{i}$$P_{i+1}$$P_{i+2}$
$P_{i-1}$$P_{i}$$P_{i+1}$$P_{i+2}$
Figure 3. A swap operation exchanging two consecutive monochromatic intervals.
The key claim is that sufficiently long alternation forces the existence of a neutral swap, i.e., a swap that does not increase $f\textsf{-width}$. To prove this, we classify intervals by their interface with the separator $X$. An $X$-ball expresses how the bags induced along an interval intersect $X$. Manageability implies that, for fixed $|X|$, $f\textsf{-width}(H)$, $\mathit{rank}(H)$, and $\Delta(H)$, only boundedly many different $X$-balls can occur (Section˜A.1). In particular, this follows from the bounded size property limiting the size of the interface, and invariant locality limiting the number of isomorphism types. Hence, if a spine contains many intervals, then many of them share the same $X$-ball. On such a homogeneous region, the effect of swapping depends only on bounded local information; a pigeonhole argument yields two consecutive intervals whose swap is neutral (Section˜A.2).
Finally, we choose among all reduced optimal forests one that minimizes an appropriate potential: first the total number of monochromatic intervals, and then the induced factorisation complexity. A neutral swap strictly decreases this potential, so a minimum contains no long alternation across any such separator. This bounded alternation translates into bounded maximal factorisations of target sets and bounded irregularity, giving the desired bounds on $\mathit{split}({\bf t},F)$ and $\mathit{mir}({\bf t},F)$.
The full proof of the lemma is the content of Appendix˜A. ∎
The combinatorial insight from Lemma˜7 may have implications beyond the algorithm presented here. It is discussed in (Bojańczyk and Pilipczuk, 2022) that the Dealteration Lemma for treewidth in a sense reflects the combinatorial core that underlies the classic FPT algorithm for checking treewidth by Bodlaender and Kloks (Bodlaender and Kloks, 1996b). This is of particular note as no analogous algorithm is known for hypergraph width measures. For our main theorem we proceed similar to (Bojańczyk and Pilipczuk, 2022), by abstractly using the combinatorial result to express our problem in terms of bounded MSO transductions. However, Lemma˜7 suggests that building on our introduction of $\mathit{split}$ and $\mathit{mir}$, and their bounded character, may form the basis of future work into a more immediate (Bodlaender-Kloks style) FPT algorithm for deciding $f\textsf{-width}$.
3.3. Efficient Construction of Transductions
The second major technical challenge in proving Theorem˜3 is the construction of a small transduction that relates tree decompositions to the optimal elimination forests of Lemma˜7.
For a tree decomposition $(T,\mathsf{bag})$ of a hypergraph $H$, $\mathit{mn}$ is a function from $V(H)$ to $V(T)$ where $u$ is mapped to the (unique) node $x$ with $u\in mrg(x)$, we say $x$ is the the margin node of $u$. We will primarily be interested in the paths in the tree decomposition induced by two margin nodes. For nodes $u,v$ of $T$, let $\mathit{Path}_{\mathit{\mathit{mn}}}(u,v)$ be the set of vertices on the shortest path from $\mathit{mn}(u)$ to $\mathit{mn}(v)$ in $T$.
Definition 8.
Let ${\bf t},F$ be a tf-pair of $H$ and let $u\in V(H)$. The stain of $u$ (w.r.t. ${\bf t},F$), denoted by $\mathit{Stain}(u,({\bf t},F))$, is the set consisting of $\mathit{mn}_{\bf t}(u)$ and the vertices of $\mathit{Path}_{\mathit{\mathit{mn}}}(u,v)$ for each child $v$ of $u$ in $F$. That is,
$\mathit{Stain}(u,({\bf t},F))=\{\mathit{mn}(u)\}\cup\bigcup_{v\in\mathit{Child}_{F}(u)}\mathit{Path}_{\mathit{\mathit{mn}}}(u,v)$
Note that Definition 8 is different from the definitions of irregularity and split above in that it considers a part of $F$ in the context of $\bf t$} rather than a part of $\bf t$} in the context of $F$. Intuitively, the stain of a vertex $u$ represents subtree of $\bf t$ that interacts with the part $F$ rooted at $u$. For the transduction it will important that the way these subtrees are coupled with each other is controlled. We capture this “coupling” more formally in terms of the conflict graph of a $\mathit{TF}$-pair ${\bf t},F$, denoted by $\mathsf{CG}({\bf t},F)$, which has $V(H)$ as the set of vertices and $u$ is adjacent to $v$ if and only if their stains intersect. Analogous to (Bojańczyk and Pilipczuk, 2022), we show that there is an MSO transduction that relates $\bf t$} to $F$ for every $\mathit{TF}$ pair, assuming that the transduction was constructed with knowledge of the chromatic number of $\mathsf{CG}({\bf t},F)$.
Lemma 9.
There is an algorithm whose input is $q>0$ and the output is a transduction $\bf Tr$} of size $O(q)$ whose input signature is $\tau_{td}$ and the output signature is $\tau_{ef}$. Let ${\bf t},F$ be a $\mathit{TF}$-pair of hypergraph $H$ and let $M(H,{\bf t})$ be the relational structure over $\tau_{td}$ induced by $H$ and $\bf t$. Then the following is true regarding ${\bf Tr}(M(H,{\bf t}))$.
  1. (1)
    Let $M^{*}\in{\bf Tr}(M(H,{\bf t}))$. Then $H(M^{*})=H$ and $F(M^{*})$ is an elimination forest of $H$.
  2. (2)
    If the chromatic number of $\mathsf{CG}({\bf t},F)$ is at most $q$, then there is a $M^{*}\in{\bf Tr}(M(H,{\bf t}))$ such that $F(M^{*})=F$.
Fortunately, it is in fact possible to bound the chromatic number of $\mathsf{CG}$ in terms of the two central parameters of our framework, $\mathit{split}$ and $\mathit{mir}$. Meaning in the context of Lemma˜7, we can consider transductions from Lemma˜9 in size bounded in terms of $\mathsf{tw}(\bf t),f\textsf{-width}(H),\mathit{rank}(H)$ and $\Delta(H)$.
Theorem 10.
There is a monotone function $\eta$ such that the chromatic number of $\mathsf{CG}({\bf t},F)$ is at most $\eta(\mathit{split}({\bf t},F),\mathit{mir}({\bf t},F),tw({\bf t}),rank(H))$.
Proof Idea.
The graph $\mathsf{CG}({\bf t},F)$ is an intersection graph of connected subgraphs of a forest, and therefore also chordal. Chordal graphs are perfect graphs, meaning their chromatic number is the size of their largest clique. We then use the fact that $\mathsf{CG}({\bf t},F)$ also enjoys the Helly property and further observations on the “density of stains” to ultimately bound the size of the largest clique in $\mathsf{CG}({\bf t},F)$. The full proof of the theorem is the content of Appendix˜C. ∎
As a final step we need a further transduction that verifies that $f\textsf{-width}$ of the forests produced by the transduction of Lemma˜9. For treewidth, this can be done implicitly, as this corresponds to limiting the size of bags themselves. To lift the overall procedure to arbitrary manageable $f\textsf{-width}$, we add another step to the transduction that filters for those elimination forests with optimal $f\textsf{-width}$. We show that such a transduction indeed always exists, with the core properties encapsulated in the following technical lemma which is the main result of Appendix˜B.
Lemma 11.
Let $f$ be a manageable width function and let $\beta$ be a size bound for $f$ as in Definition 2. Then, for each $k\geq 0$ and positive integer $r$, the following statements hold.
  1. (i)
    There is an MSO sentence $\varphi^{*}_{f,k,r}$ over the signature $\tau_{ef}$ such that for a structure $M$ over the signature, $M\models\varphi^{*}_{f,k,r}$ if and only if $F(M)$ is of $f\textsf{-width}$ at most $k$.
  2. (ii)
    There is an algorithm that outputs $\varphi_{f,k,r}^{*}$ with input $k,r$ in time that depends only on a computable function in $k,r$ and $\beta$.
  3. (iii)
    Let $\mathcal{H}_{k,r}$ be all hypergraphs with $\mathit{rank}(H)\leq r$ and $f_{H}(V(H)\leq k$. Then there is an algorithm that computes the set $V=\{f_{H}(V(H))\mid H\in\mathcal{H}_{k,r}\}$ in time depending on $k,r$, and $\beta$. In consequence, $|V|$ can also be bounded in terms of $k,r,\beta$.
3.4. Proof of Theorem 3
Putting together the contents of the previous two sections forms the spine of our algorithm. We will show that there is an MSO transduction that represents a sound and complete algorithm for computing an optimal width elimination forest from an initial tree decomposition. Using the following result of Bojańczyk and Pilipczuk and the bounds we obtain on the size of our transduction we then observe that the resulting algorithm is indeed fixed-parameter tractable.
Proposition 12 (Theorem 6.1 in (Bojańczyk and Pilipczuk, 2022)).
[…]
$f(\|\mathbf{Tr}\|,\mathsf{tw}(\mathbb{A}))\ (\|\mathbb{A}\|+m),$
where $m$ is the size of the output structure (or 0 if empty) and $f$ is computable.
Proof of Theorem 3.
The algorithm first computes $p=\beta(k,rank(H))$ and runs an off-the-shelf algorithm for exact treewidth computation parameterized by p1p-1. We note that, due to the bounded size property, if $f\textsf{-width}(H)\leq k$ then $\mathsf{tw}(H)\leq p-1$. Therefore, if the treewidth computation algorithm rejects, our algorithm also rejects.
Otherwise, we have a tree decomposition ${\bf t}=(T,B)$ of $H$ with treewidth at most p1p-1. We then compute $p_{1}=\gamma(tw({\bf t}),k,rank(H),\Delta(H))$ with $\gamma$ as in Lemma 7, and $q=\eta(p_{1},p_{1},tw({\bf t}),rank(H))$ with $\eta$ as in Theorem 10. Finally, we apply the algorithm as in Lemma 9 with input $q$ to create a transduction $\bf Tr$} as specified in the lemma.
Claim 13.
Assume that $f\textsf{-width}(H)\leq k$.
Then there is $M^{*}\in{\bf Tr}(M(H,{\bf t}))$ such that $f\textsf{-width}(F(M^{*}))=f\textsf{-width}(H)$.
Proof of Claim. By Lemma 7 and the monotonicity of function $\gamma$ as in the lemma, there is an elimination forest $F$ of $H$ with $f\textsf{-width}(F)=f\textsf{-width}(H)$ and such that both $\mathit{split}({\bf t},F)$ and $mit({\bf t},F)$ are at most $p_{0}$. Next, by Theorem 10 and monotonicity of $\eta$, the chromatic number of $\mathsf{CG}({\bf t},F)$ is at most $q$. It follows from Lemma 9 that $Tr(M(H,{\bf t}))$ contains a model $M_{1}$ with $F(M_{1})=F$. ∎
Next, we apply the algorithm as in the last item of Lemma 11 to produce the set $\mathit{Val}_{f,k,r}$ of values as specified in Item˜(iii) of the lemma.
Our algorithm explores the values $a\in\mathit{Val}_{f,k,r}$ in the increasing order. For each value $a$, a subprocedure tests whether $f\textsf{-width}(H)\leq a$ and either returns a corresponding tree decomposition or rejects. In the latter case it is guaranteed that $f\textsf{-width}(H)>a$. If the subprocedure rejects for all $a\in\mathit{Val}_{f,k,t}$, then by definition of $\mathit{Val}_{f,k,r}$ the overall algorithm is safe to reject too. Otherwise, let $a$ be the first value for which a tree decomposition is returned. Once again, according to the properties of $\mathit{Val}_{f,k,r}$, $f\textsf{-width}(H)=a$ and hence the algorithm is safe to return the obtained tree decomposition. It thus remains to describe the subprocedure for testing $f\textsf{-width}(H)\leq a$ for a specific value $a\leq k$.
By Lemma 11, $\varphi^{*}_{f,a,r}$ (as in the description of the lemma) can be designed by an algorithm in at most $g_{0}(k,r)$ time. We extend the transduction $\bf Tr$} by appending one domain filtering transduction with $\varphi^{*}_{f,a,r}$ being the underlying MSO formula. We call the resulting transduction $\bf Tr^{*}_{\mathit{a}}$}.
Claim 14.
The following two statements hold.
  1. (1)
    If $f\textsf{-width}(H)\leq a$ then ${\bf Tr}^{*}_{a}(M(H,{\bf t}))\neq\emptyset$. Moreover, for each $M^{*}\in{\bf Tr}^{*}_{a}(M(H,{\bf t}))$, $H(M^{*})=H$ and the $f$-width of $F(M^{*})$ is at most $a$.
  2. (2)
    If $f\textsf{-width}(H)>a$ then ${\bf Tr}^{*}_{a}(M(H,{\bf t}))=\emptyset$.
Proof of Claim. Assume that $f\textsf{-width}(H)\leq a$. In particular, this means that $f\textsf{-width}(H)\leq k$. Hence, we apply Claim 13 and conclude that the output of ${\bf Tr}(M(H,{\bf t}))$ contains at least one model $M^{*}$ as specified in the statement of this claim. This model satisfies the underlying MSO of the last elementary transduction of ${\bf Tr}^{*}_{a}(M(H,{\bf t}))$ and hence the latter is not empty. Let $M^{*}\in{\bf Tr}^{*}_{a}(M(H,{\bf t}))$. This means that $M^{*}\in{\bf Tr}(M(H,{\bf t}))$. Hence, by Lemma 9, $H(M^{*})=H$ and $F(M^{*})$ is an elimination forest for $H$. This means that $M^{*}$ passes through the filtering of the last stage of ${\bf Tr}^{*}_{a}(M(H,{\bf t}))$ and thus $f\textsf{-width}(F)\leq a$. This proves the first statement of the claim.
For the second statement, assume that $f\textsf{-width}(H)>a$ and yet ${\bf Tr^{*}}(M(H,{\bf t}))\neq\emptyset$. Let $M^{*}\in{\bf Tr}^{*}_{a}(M(H,{\bf t}))$. This means that $M^{*}\in{\bf Tr}(M(H,{\bf t})$. By Lemma 9 and the assumption, $f\textsf{-width}(F(M^{*}))>a$. This means that $M^{*}$ does not pass through the filtering of the last step of ${\bf Tr^{*}}_{a}(M(H,{\bf t}))$ and hence cannot belong to the output. This contradiction proves the second statement of the claim. ∎
Having constructed ${\bf Tr}^{*}_{a}$, we feed it along with $M(H,{\bf t})$ to the input of the algorithm $\mathcal{A}$ as in Proposition˜12. According to the theorem, $\mathcal{A}({\bf Tr^{*}}_{a},M(H,{\bf t}))$ returns an elimination forest of $H$ of width at most $a$ or rejects and, in the latter case, it is guaranteed that $f\textsf{-width}(H)>a$. If an elimination forest returned it only remains to turn it into a tree decomposition of $H$ induced by $F$ that can be done in a polynomial time. It thus only remains to verify the FPT time bound for $\mathcal{A}({\bf Tr}^{*}_{a},M(H,{\bf t}))$.
The runtime of $\mathcal{A}({\bf Tr}^{*}_{a},M_{0})$ is FPT parameterized by the size of ${\bf Tr^{*}}_{a}$ and tw(M(H,t))tw(M(H,{t})). By our initial choice of $p$, we have that tw(M(H,t))tw(M(H,{t})) is at most $\beta(k,rank(H))$. The size of ${\bf Tr}^{*}_{a}$ is the size of $\bf Tr$} plus the size of $varphi^{*}_{f,a,r}$. The size of $\bf Tr$} us upper bounded by a function of $q$ which itself is upper bounded by a function of $k,rank(H),\Delta(H)$ by definition. Finally, the length of the last step is upper bounded by a function of $k$ and $rank(H)$ by Lemma 11. ∎
While our proof does not directly provide an easily implementable algorithm, it does finally confirm that FPT checking of $\mathsf{ghw}$ and $\mathsf{fhw}$, and in fact computation of witnessing decompositions, is possible. We believe that the key insights for obtaining Theorem˜3, as well as establishing that FPT algorithms are theoretically possible, can guide the development for practical algorithms in future work. While parameters $f\textsf{-width}(H)$ and $\mathit{rank}(H)$ are central to our approach, the degree is only necessary to obtain the bound in Lemma˜7. Its role is an interesting topic for future work, and we discuss the matter in detail in Section˜5.
4. Computation and Approximation of Adaptive Width Variants
Let $\mathcal{F}$ be a family of width functions. The $\mathcal{F}$-width of hypergraph $H$ is defined as the $\sup_{f\in\mathcal{F}}f\textsf{-width}(H)$. Width measures based on such a whole class of width functions have become a central element of the study of the complexity of database query answering (Marx, 2011, 2013). Despite their importance, we are not aware of any algorithmic study of the complexity of deciding such width measures. Here we show how our framework for manageable width functions also applies to some of these cases.
Any function $f:V(H)\to\mathbb{R}_{\geq 0}$ naturally extends to a width function by setting $f(U)=\sum_{v\in U}f(v)$. Such width functions are called modular. Let us additionally require that $f$ is a fractional independent set, i.e., $f(e)\leq 1$. Let $\mathcal{F}$ be the class of all such modular width function induced by fractional independent sets (for a given $H$). The $\mathcal{F}$-width is then the so-called adaptive width of $H$, a parameter that characterizes when deciding conjunctive queries is in FPT (Marx, 2013). For simplicity we fix $\mathcal{F}$ to this class of modular width functions induced by fractional independent sets for the rest of this paper.
In this section, we discuss how algorithms for two restricted versions of adaptive width follow from the framework for FPT $f\textsf{-width}$ checking described above.
Definition 1.
Let $H$ be a hypergraph and $\delta\geq rank(H)$ be an integer. We define $\mathcal{F}_{\delta}$ as the family of all modular width functions induced by fractional independent sets $f$ where each $v\in V(H)$, $f(v)=\frac{a}{\delta}$ where $a\in\mathbb{N}_{0}$ and $0\leq a\leq\delta$. Furthermore, define $\mathcal{F}^{*}_{\delta}$ as the set of all modular width functions induced by fractional independent sets $f$ such that for each $v\in V(H)$ either $f(v)=0$ or $f(v)\geq\frac{1}{\delta}$.
Thus, $\mathcal{F}_{\delta}$-width is the discretized version of adaptive width with bounded precision expressed by $\delta$. $\mathcal{F}^{*}_{\delta}$-width relaxes this restriction further and only excludes functions that assign very small non zero weights. To apply our framework, first observe that any fixed $f\in\mathcal{F}_{\delta}$ satisfies the weak monotone separability and automatizability properties of Definition˜2. The bounded size property holds on the $H^{f+}$, the induced subhypergraph on vertices where $f(v)\neq 0$. Since $f\textsf{-width}(H)=f\textsf{-width}(H^{f+})$, simply operating on $H^{f+}$ will be sufficient for us. The only missing part is invariant locality. Intuitively, we fix each $f\in\mathcal{F}_{\delta}$ as vertex labels on the hypergraph. On a technical level, this yields a labelled hypergraph (each node labelled with its value under $f$) for which $f$ indeed satisfies invariant locality. That is, at the technical level where we use our previous machinery on the width functions from $\mathcal{F}_{\delta}$, they are indeed manageable, and the machinery discussed in the previous section works without change.
The main result of this section is a reformulation of Theorem 2 as stated below.
Theorem 2.
[…]
We need to introduce some further notation. Let $\delta\in\mathbb{N}$ and let $\mathcal{U}=\{U_{0},\dots,U_{\delta}\}$ be a weak partition of $V(H)$. We write $f[U_{0},\dots,U_{\delta}]$ (or simply $f[\mathcal{U}]$) for the modular width function that, for every vertex $v$, assigns $f(v)=\frac{a}{\delta}$ when $v\in U_{a}$. For clarity we write $\hat{f}[\mathcal{U}]$ for the modular width function induced by the vertex weight function $f[\mathcal{U}]$.
The central statement for the proof of Theorem 2 is to is the following theorem. It follows from the same underlying machinery as Theorem˜3, but instead of turning the width check transduction directly into an algorithm, we translate it to an MSO formula where a modular width function is given as input via second-order variables.
Theorem 3.
There is an algorithm $\mathcal{A}(H,{\bf t},k,\delta)$, where $H$ is a hypergraph, $\bf t$} is a tree decomposition of $H$, $k$ is a non-negative rational, and $\delta\in\mathbb{N}$, that returns an MSO formula $\varphi(U_{0},\dots,U_{\delta})$ where $\mathcal{U}=\{U_{0},\dots,U_{\delta}\}$ are free (monadic) second-order variables such that: $M\not\models\varphi(U_{0},\dots,U_{\delta})$ exactly if $\mathcal{U}$ is a weak partition of $V(H)$, $f[\mathcal{U}]$ is a fractional independent set, and $\hat{f}[\mathcal{U}]$-width(H)>k(H)>k.
The runtime of the algorithm is FPT parameterized by the treewidth of $\bf t$}, $k$, $rank(H)$, $\Delta(H)$ and $\delta$ and the size of the resulting formula is upper-bounded by a function depending on these parameters.
To prove Theorem˜2 we will extend the formula from Theorem˜3 by (indirect) second-order quantification over the possible width functions. In particular,
with the notation as in Theorem˜3, let $\psi=\forall U_{0}\dots\forall U_{\delta}\ \varphi(U_{0},\dots,U_{\delta})$.
Claim 4.
$M\models\psi$ if and only if $\mathcal{F}_{\delta}$-width of $H$ is at most $k$.
Proof of Claim. Assume first that $M\models\psi$. For the sake of contradiction, assume that the $\mathcal{F}_{\delta}$-width of $H$ is greater than $k$. This means that there is a function $f\in\mathcal{F}_{\delta}$ such that the $f\textsf{-width}$ of $H$ is greater than $k$. By definition of $\mathcal{F}_{\delta}$, there is a weak partition $U^{\prime}_{0}\dots U^{\prime}_{\delta}$ such that for each $i\in\{0,\dots\delta\}$, $U^{\prime}_{i}$ consists of exactly those elements $v$ such that $f(v)=\frac{i}{\delta}$. Clearly, $f=f[U^{\prime}_{0},\dots U^{\prime}_{\delta}]$ is a fractional independent set. By definition $M\not\models\varphi(U^{\prime}_{0},\dots U^{\prime}_{\delta})$. Then $M\not\models\psi$ as witnessed by $U_{0}=U^{\prime}_{0},\dots,U_{\delta}=U^{\prime}_{\delta}$ thus contradicting our assumption.
Assume now that $M\not\models\psi$. This means that there are $U_{0}=U^{\prime}_{0},\dots U_{\delta}=U^{\prime}_{\delta}$ such that $M\not\models\varphi(U^{\prime}_{0}\dots,U^{\prime}_{\delta})$. By definition, this means that $f=f[U^{\prime}_{0},\dots,U^{\prime}_{\delta}]$ is a fractional independent set of $H$ and the $f\textsf{-width}$ of $H$ is greater than $k$. By definition of $f$, clearly $f\in\mathcal{F}_{\delta}$. It follows that $\mathcal{F}_{\delta}$-width of $H$ is greater than $k$. ∎
We now sketch the design of the required algorithm. We observe first that $tw(H)$ is at most twice $\mathcal{F}_{\delta}$-width multiplied by $rank(H)$. The first step of the algorithm is running a linear time algorithm for treewidth computation parameterized by $2\cdot k\cdot rank(H)$. If the treewidth construction algorithm rejects then our algorithm can safely return $False$. Otherwise, the treewidth construction algorithm returns a tree decomposition ${\bf t}=(T,{\bf B})$ of treewidth at most $k\cdot rank(H)\cdot\Delta(H)$. We run $\mathcal{A}(H,{\bf t},k,\delta)$ as in the statement of Theorem 3. We note that, by definition of $\bf t$}, the runtime of the algorithm $\mathcal{A}$ is FPT parameterized by $k,rank(H),\Delta(H)$, and $\delta$. It follows that the size of $\varphi(U_{0},\dots,U_{\delta})$ is upper bounded by such a function of these parameters. Clearly, that the runtime of construction of $\psi$ is FPT parameterized by the same parameters and the size of $\psi$ can be upper bounded by a function dependent on $k,rank(H),\Delta(H)$, and $\delta$.
By Claim 4, it only remains to test whether $M(H,{\bf t})\models\psi$. This can be done within the required runtime according to Courcelle’s theorem (Courcelle and Engelfriet, 2012). The full proof details for Theorem˜2 are presented in Appendix˜E.
As a corollary of Theorem 2 we can show that $\mathcal{F}^{*}_{\delta}$-width is FPT approximable with ratio $2$ through a rounding argument that connects it to $\mathcal{F}_{\delta}$-width. We then use the algorithm of Theorem 2 with the width parameter being $k$.
Corollary 5.
[…]
It follows from Corollary 5, that if there is $\delta=\eta(k,rank(H),\Delta(H))$ such that for all $H$, $\mathcal{F}^{*}_{\delta}=\mathcal{F}$ then adaptive width has a ratio $2$ FPT approximation parameterized by the width $k$, $rank(H)$, and $\Delta(H)$. Thus existence of such a function $\eta$ is an interesting open question for future work.
5. Conclusion & Future Work
In this paper we have developed a general framework for fixed-parameter tractable computation of hypergraph width measures. By introducing the class of manageable width functions and as a special case we obtain the first exact FPT algorithms for deciding $\mathrm{ghw}(H)\leq k$ and $\mathrm{fhw}(H)\leq k$ under any non-trivial parameterisation. Our framework further yields an exact FPT algorithm for discretized adaptive width, as well as a factor-$2$ FPT approximation algorithm for the relaxed variant $F^{*}_{\delta}$-width.
A natural question for further research is whether the parameter $\Delta(H)$ is necessary for fixed-parameter tractability of $f\textsf{-width}$ computation. For the related discussion, we find it easier to confine ourselves to a specific parameter $\mathsf{ghw}$. Then the open question is as follows: is testing $\mathsf{ghw}(H)\leq k$ fixed-parameter tractable if parameterized by $k$ and $\mathit{rank}(H)$?
In order to upgrade the machinery of this paper to the parametrization by $k$ and $\mathit{rank}(H)$ only, it is enough to get rid of $\Delta(H)$ in the statement of Lemma 8. Formally this leads to the following question.
Open Question 1.
Is there a function $h$ such that the following holds. Let $F$ be an elimination forest of $H$ and let $(Red,X,Blue)$ be a red-blue partition of $V(H)$. Let $B$ be a context factor of $F$ and suppose that the number of monochromatic intervals of the spine of $B$ is at least $h(|X|,\mathsf{ghw}(F),rank(H))$. Then $B$ admits a neutral swap (cf., Lemma˜8).
A positive answer to Question 1 will lead to an immediate upgrade of the algorithm in this paper to parameterization by $k$ and $rank(H)$. In order to answer the question negatively, a counterexample of the following kind is needed. Pick $k,r>2$ and an infinite set of hypergraphs $H_{1},H_{2},\dots$, their respective elimination forests $F_{1},F_{2}\dots$ and respective context factors $B_{1},B_{2},\dots$ with the number of monochromatic intervals of these factors monotonically increasing and yet none of $B_{i}$ admitting a neutral swap.
We conclude the section with an open question related to adaptive width.
Open Question 2.
[…]
A positive answer to Question 2, combined with the results of the previous section will imply a constant ratio – in particular ratio $2c$ – FPT approximation algorithm for adaptive width parameterized by the width, the rank, and the max-degree. Once again, the question is of a purely combinatorial nature.
Appendix A Proof of the Reduced Hypergraph Dealteration Lemma (Lemma 7)
In this section we prove the Reduced Dealteration Lemma.
We start from introducing new terminology.
Definition 1.
An elimination forest $F$ of a hypergraph $H$ is reduced if for each non-leaf vertex $u$ and each child $v$ of $u$, $u$ is adjacent in $H$ to $F_{v}$.
The essential property of reduced elimination forests is stated in the following.
Lemma 2.
Let $H$ be a hypergraph and let $F$ be a reduced elimination forest for $H$. Then, for each $u\in V(H)$, $H[V(F_{u})]$ is connected. Moreover, for each non-leaf vertex $u$ and each child $v$ of $u$, $H[F_{v}\cup\{u\}]$ is also connected.
Proof.
By bottom up induction on $F$. If $u$ is a leaf then the statement is obvious. Otherwise, the tree factor of each child $v$ of $u$ is a connected subset and $u$ itself is connected to each $F_{v}$ by definition of a reduced forest. ∎
In what follows, we fix a partition $(Red,X,Blue)$ of $V(H)$ so that $H$ has no hyperedge $e$ containing both a red and a blue vertex. We refer to the vertices of $Red$ as red vertices and to the vertices of $Blue$ as blue vertices. We say that a set $U\subseteq Red\cup Blue$ is monochromatic if either $U\subseteq Red$ or $U\subseteq Blue$. The reasoning provided below will be mostly symmetric for colours. For $u\in Red\cup Blue$, an important part of such a reasoning is which of $Red,Blue$ is of the same colour as $u$ and which of the opposite one. For this purpose, we introduce the notation $Native(u)$ and $Foreign(u)$. In particular, if $u\in Red$ then $Native(u)=Red$ and $Foreign(u)=Blue$. If, on the other hand, $u\in Blue$ then $Native(u)=Blue$ and $Foreign(u)=Red$.
Lemma 3.
Let $H$ be a hypergraph and let $F$ be a reduced elimination forest for $H$. Let $U\subseteq Red\cup Blue$. If there is $u\in V(H)$ such that $U=F_{u}$ then $U$ is monochromatic. Furthermore, if there is $u\in V(H)$ and a child $v$ of $U$ such that $U=F_{v}\cup\{u\}$ then $U$ is also monochromatic.
Proof.
By Lemma 2, $U$ is connected. On the other hand, there is no edge connecting $U\cap Red$ and $U\cap Blue$. Hence, one of these sets must be empty. ∎
In order to proceed, we need to recall, for a constant factor $B$ of $F$, the notions of the root $rt(B)$ of $B$, the parent $pr(B)$ of the appendices of $B$ and the spine $spine(B)$. Let us note that a context factor is uniquely determined by its spine and the appendices. Indeed, let $P$ be a directed path of $F$ whose final vertex is not a leaf. Let $W$ be a non-empty subset of children of the final vertex of $P$. Then the context factor $Cont_{F}(P,W)$ is obtained by adding to $P$ each $V(F_{u})$ such that $u$ is either a child of a non-final vertex of $P$ other than its successor on $P$ or a a child of the final vertex of $P$ that does not belong to $W$.
Lemma 4.
Suppose that $F$ is reduced. Let $P$ be a directed path of $F$ whose final vertex is not a leaf. Let $W$ be a non-empty subset of children of the final vertex of $P$. Suppose that $Cont_{F}(P,W)\subseteq Red\cup Blue$ and $V(P)$ is monochromatic. Then $Cont_{F}(P,W)$ is monochromatic.
Proof.
It is enough to show that for each $u\in V(P)$ and for each child $v$ of $u$ such that $V(F_{v})\subseteq Cont_{F}(P,W)$, $F_{v}\cup\{u\}$. Since, by definition, $F_{v}\cup\{u\}\subseteq Red\cup Blue$, this is immediate from Lemma 3. ∎
Definition 5.
Let $U\subseteq Red\cup Blue$ be a context factor of an elimination forest $F$ of a hypergraph $H$ Let $P$ and $W$ be the spine and the appendices of $U$ respectively . We call a maximal monochromatic subpath of $P$ a (colour) interval. In other words $P=P_{1}+\dots+P_{q}$ where $P_{1},\dots,P_{q}$ are the intervals of $P$. For each $1\leq i\leq q$, we introduce a set $W_{i}$ of local appendices as follows. $W_{q}=W$ and for each $1\leq i\leq q-1$, $W_{i}$ is a singleton containing the first vertex of $P_{i+1}$, We call the factors $Cont_{F}(P_{i},W_{i})$ for $1\leq i\leq q$ the interval factors of $U$.
Corollary 6.
With the notation as in Definition 5, and, under assumption that $F$ is reduced, each interval factor of $U$ is monochromatic.
Proof.
Immediate from Lemma 4. ∎
We are now going to formally introduce the notion of a swap: a modification of $F$ that swaps two interval factors of $U$ and hence reduces the number of factors by two. We will next define a neutral swap as one that does not increase th $f$-width of the resulting forest. The main engine of the Reduced Dealteration Lemma is proving that if $F$ is reduced and $U$ contains many interval factors then $U$ admits a neutral swap. In this section we will prove the statement of the main engine. We postpone to Section A.3 the description of the way it is harnessed to prove the Reduced Dealteration Lemma.
Definition 7.
Continuing on Definition 5, assume that $q\geq 4$. Let 1<i<q11<i<q-1. Let $(u_{1},u_{2})$ be the edge from the last vertex of $P_{i-1}$ to the first vertex of $P_{i}$. Let $(v_{1},v_{2})$ be the edge from the last vertex of $P_{i}$ to the first vertex of $P_{i+1}$. Finally, let $(w_{1},w_{2})$ be the edge connecting the last vertex of $P_{i+1}$ and the first vertex of $P_{i+2}$. Let $F^{\prime}$ be obtained from $F$ by removal of edges $(u_{1},u_{2})$ , $(v_{1},v_{2})$, and $(w_{1},w_{2})$ and adding edges $(u_{1},v_{2})$, $(w_{1},u_{2})$, $(v_{1},w_{2})$. We call $F^{\prime}$ a swap of $F$ and the intervals being swapped are $P_{i}$ and $P_{i+1}$.
We say that the swap is neutral if $f\textsf{-width}(F^{\prime})\leq f\textsf{-width}(F)$. In particular, in this case that $U$ admits a neutral swap (of intervals $P_{i}$ and $P_{i+1}$).
$P_{i-1}$$P_{i}$$P_{i+1}$$P_{i+2}$$P_{i-1}$$P_{i}$$P_{i+1}$$P_{i+2}$
Figure 4. A swap $F^{\prime}$ of $F$.
Lemma 8.
There is a monotone function $\kappa$ such that the following holds. Let $F$ be a reduced elimination forest of $H$ of minimal width (i.e., $f\textsf{-width}(F)=f\textsf{-width}(H)$), and let $U$ be a coloured context factor. If the spine of $U$ has more than $\kappa\bigl(|X|,f\textsf{-width}(H),\mathit{rank}(H),\Delta(H)\bigr)$ monochromatic intervals, then $U$ admits a neutral swap.
See 8
We now proceed proving Lemma 8. In what follows throughout the proof, the notation is as in the lemma, unless otherwise explicitly specified.
First of all, we extend our terminology about bags.
Definition 9 ($X$-bags, native bags, and foregin bags).
Let $u\in V(H)$. We refer to $bag_{F}(u)\cap X$ as the $X$-bag of $u$ and denote it as Xbag(u)X-bag(u). For $u\in Red\cup Blue$, we refer to $bag_{F}(u)\cap Native(u)$ as the native bag of $u$ denoted by $nbag_{F}(u)$ and to $bag_{F}(u)\cap Foreign(u)$ as the foreign bag of $u$ denoted by $fbag_{F}(u)$.
Definition 10.
Let $W$ be a connected component of $H[bag_{F}(u)]$. We say that $W$ is an $X$-component if $W\cap X\neq\emptyset$. Note that otherwise $W$ is monochromatic as no hyperedge of $H$ connects both a red and a blue vertex. If $W$ is of the same colour as $u$ then we say that $W$ is a native component of $u$. Otherwise, $W$ is a foreign component of $u$.
Based on the last definition we partition $bag_{F}(u)$ into unions of connected components subject to their classification.
Definition 11.
We call the union of all $X$-components of $H[bag_{F}(u)]$ the $X$-ball of $u$ (w.r.t. $F$) and denote it by $X-ball_{F}(u)$. Further on, we refer to the union of all native components of $H[bag_{F}(u)]$ as the native ball of $u$ (w.r.t. $F$) and denote it by $nball_{F}(u)$. Finally, we refer to the union of foreign components of $bag_{F}(u)$ as the foreign ball of $u$ (w.r.t. $F$) and denote it by $fball_{F}(u)$.
For the reasoning that follows, it is important that the $X$-balls, native balls and foreign balls of a bag are edge disjoint. It thus follows from the monotone weak separability of $f$ that
Proposition 12.
For each $u\in V(H)$, $f(bag_{F}(u)=g_{f}(f(X-ball(u)),f(nball_{F}(u)),f(fball_{F}(u)))$.
Definition 13.
With the notation as in Definition 5, let $P^{\prime}$ be a maximal subpath of such that for each $u\in V(P)$, $X-ball_{F}(u)$ is the same. We refer to $P^{\prime}$ an $X$-interval of $U$.
Note that $X$-intervals demonstrate a partition of $P$ alternative to the colour interval. In particular, it is always guaranteed that the number of $X$-intervals is bounded.
Lemma 14.
The number of $X$-intervals of $U$ is at most $\kappa_{0}(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$ where $\kappa_{0}$ is a monotone function.
The proof of Lemma 14 is provided is Subsection A.1. The critical role of the lemma is that it is the only statement in the whole proof of our result where the bounded degree is essential.
Lemma 14 in fact demonstrates that $U$ can be represented as the disjoint union of a bounded number of $X$-homogenous context factors, the corresponding notion is defined below.
Definition 15.
Let $U_{0}\subseteq Red\cup Blue$ be a context factor and let $P_{0}$ be the spine of $U_{0}$. We say that $U_{0}$ is $X$-homogenous if all the vertices of $P_{0}$ have the same $X$-ball.
Lemma 16.
There is a function $\kappa_{1}$ such that the following holds. Let $U_{0}$ be an $X$-homogenous context factor with spine $P_{0}$ that has at least $\kappa_{1}(f\textsf{-width}(H),rank(H))$ colour factors. Then $U_{0}$ admits a neutral swap of two non-terminal (neither first nor last) intervals
The proof of Lemma 16 is provided in Subsection A.2. Now, we are ready to prove Lemma 8.
Proof of Lemma˜8.
Let $r$ be the number of colour intervals of $P$. Assume that $r\geq\kappa_{1}(f\textsf{-width}(H),rank(H))\cdot\kappa_{0}(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$ where $\kappa_{0}$ is as in Lemma 14 and $\kappa_{1}$ is as in Lemma 16. We prove that in this case $U$ admits a neutral swap. Let $P^{\prime}_{1},\dots,P^{\prime}_{q}$ be the partition of $P$ into $X$-intervals. According to Lemma 14, $q\leq\kappa_{0}(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$. It is not hard to see that $U$ is the disjoint union of context factors $U_{1},\dots,U_{q}$ with $P^{\prime}_{i}$ and $W^{\prime}_{i}$ be the respective spine and appendices for each $U_{i}$ where $W^{\prime}_{i}$ is a singleton consisting of the first vertex of $P^{\prime}_{i+1}$ if $i<q$ and $W^{\prime}_{q}=W$. For $1\leq i\leq q$, let $r_{i}$ be the number of colour intervals of $U_{i}$. Clearly, $r\leq r_{1}+\dots r_{q}$ as each colour interval of each $U_{i}$ is a subset of exactly one colour interval $U$. It thus follows from the pigeonhole principle that some $U_{i}$ has at least $\kappa_{1}(f\textsf{-width}(H),rank(H))$ colour intervals. By Lemma 16 and the definition of a neutral swap, $U_{i}$ admits a swap of two non-terminal colour intervals that does not increase the $f\textsf{-width}$ of the resulting elimination forest $F^{\prime}$. We note that each non-terminal colour interval of $U_{i}$ is also a colour interval of $U$ and two such consecutive intervals are also consecutive for $U$. We conclude that their swap is also a swap for $U$ that does not increase the $f\textsf{-width}$ of $F^{\prime}$ and hence a neutral swap of $U$ as required. ∎
A.1. Bounding the Number of $X$-Intervals (Lemma 14)
Let $\beta=\beta[f]$ be as in the bounded size property of manageable functions and let us define for this section $p:=\beta(f\textsf{-width}(H),rank(H))$.
Lemma 17.
For each $u\in V(P)$, each vertex $w\in X-ball_{F}(u)$ is at distance at most p1p-1 from some vertex of $X$.
Proof.
Indeed, let $w\in X-ball_{F}(u)$. Then there is a connected component $W$ of $H[bag_{F}(u)]$ such that $W\cap X\neq\emptyset$ and $w\in W$. It follows that there is a path $Q$ between $w$ and some vertex $x\in X$ such that $V(Q)\subseteq bag_{F}(u)$. It follows from the bounded size condition that $p\geq|bag_{F}(u)|\geq|V(Q)|$. In other words $Q$ contains at most p1p-1 edges and hence $w$ is at distance at most p1p-1 from $X$ as required. ∎
Lemma 18.
Let $\bf Y$} be the set of all $X$-balls of the vertices of $P$. Then $|{\bf Y}|\leq(|X|\cdot\sum_{i=0}^{p-1}\Delta(H)^{p-1})^{p}$.
Proof.
Let $W(X)$ be the set of vertices of $H$ that are at distance at at most p1p-1 from $X$. It is not hard to see that $|W(X)|\leq|X|\cdot\sum_{i=0}^{p-1}\Delta(H)^{p-1}$. On the other hand, it follows from the combination of Lemma 17 and the bounded size condition that each element of $\bf Y$} is a subset of $W(X)$ of size at most $p$ thus implying the desired upper bound. ∎
Let $P^{\prime}_{1},\dots,P^{\prime}_{q}$ be the partition of $P$ into $X$-intervals and let $X_{1},\dots,X_{q}$ be the corresponding $X$-balls of vertices of $P^{\prime}_{1},\dots,P^{\prime}_{q}$. Had $X_{1},\dots,X_{q}$ been all distinct, Lemma 14 would have immediately followed from Lemma 18. However, this is not necessarily so as, for instance, a sequence $\{v_{1}\},\{v_{1},v_{2}\}$ and then back $\{v_{1}\}$ is perfectly possible. This matter can be addressed through the notion of peaks.
Definition 19.
We say that $X_{i}$ is a peak if $i=q$ or if $i<q$ and $X_{i}\setminus X_{i+1}\neq\emptyset$.
Lemma 20.
The peaks are all distinct.
Proof.
Indeed, assume the opposite. Then there are $1\leq i<j\leq q$ such that $X_{i}$ and $X_{j}$ are peaks and $X_{i}=X_{j}$. Let $v\in X_{i}\setminus X_{i+1}$. We note that $X_{i}\cap X_{i+1}\neq\emptyset$. Indeed, by definition $X_{i}\cap X=X_{j}\cap X\neq\emptyset$. By the connectivity property of tree decompositions, $X_{i}\cap X\subseteq X_{i+1}$. Let $Q_{v}$ be a path from $X_{i}\cap X$ to $v$ in $H[X_{i}]$.
Claim 21.
Let $w\in P^{\prime}_{i+1}$. Then $V(Q_{v})$ is not a subset of $bag_{F}(w)$.
Proof of Claim. Indeed, assume the opposite. Then in $H[bag_{F}(w)]$ $v$ is connected to a vertex of $X$ and hence belongs to $X-bag_{F}(w)=X_{i+1}$, a contradiction. ∎
Note a subtle point of Claim 21. We do not claim that $v\notin bag_{F}(w)$, we claim that at least one of $V(Q_{v})$ is outside of $bag_{F}(w)$,
It follows from the tree decomposition axioms that all the vertices of $P$ whose bags w.r.t. $F$ contain $Q_{v}$ as a subset form a subpath $P_{v}$ of $P$. By assumption both $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are subpaths of $P_{v}$. But then for each $a>0$ such $i+a\leq j$, $P^{\prime}_{i+a}$ is also a subpath of $P_{v}$. In particular, $P^{\prime}_{i+1}$ is a subpath of $P_{v}$. However, this is contradicts Claim 21
Proof of Lemma˜14.
We partition $[q]$ into maximal intervals $I_{1},\dots I_{a}$ such that each $I_{b}$ is a chain in the following sense. Let $b_{1}<b_{2}\in I_{b}$, Then $X_{b_{1}}\subseteq X_{b_{2}}$. Let $j$ be the final index of $I_{b}$. Then, due to to the maximality of $I_{b}$, $X_{j}$ is a peak. We thus conclude from the combination of Lemma 18 and Lemma 20 that $a\geq(|X|\cdot\sum_{i=0}^{p-1}\Delta(H)^{p-1})^{p}$. On the other hand, the elements of each $I_{b}$ are in a bijective correspondence with a chain of sets of size at most $p$ each. We thus conclude that $q=|I_{1}|+\dots+|I_{a}|\leq p\cdot(|X|\cdot\sum_{i=0}^{p-1}\Delta(H)^{p-1})^{p}$. as required. ∎
A.2. Neutral Swaps on Spines with Many Factors (Lemma 16)
We first prove a combinatorial statement about sequences of real numbers and then harness this statement to establish a proof of Lemma 16.
Definition 22.
Let $S=(a_{1},\dots,a_{r})$ be a sequence of numbers. The quadruple $(b_{1},b_{2},b_{3},b_{4})$ of four consecutive elements of $S$ is called mutable if $b_{3}\geq b_{1}$ and $b_{2}\geq b_{4}$.
Theorem 23.
Let $I$ be a finite set of reals and let $S=(s_{1},\dots,s_{r})$ be a sequence such that $s_{i}\in I$ for each $i\in[r]$. Suppose that $r\geq 4|I|$. Then $S$ has a mutable quadruple.
Proof.
By induction on $|I|$. If $|I|=1$ then we observe that $4$ consecutive $a,a,a,a$ form a mutable sequence.
Assume now that $|I|>1$ and let $a$ be the largest element of $I$. Consider a sequence $S$ as specified in the statement of the lemma and having length at least $4|I|$.
Suppose that $S$ has $5$ consecutive elements of the form $a_{1},a_{2},a,a_{3},a_{4}$. If $a_{2}\geq a_{3}$ then $a_{1},a_{2},a,a_{3}$ is a mutable quadruple. Otherwise, if $a_{2}<a_{3}$ then $a_{2},a,a_{3},a_{4}$ is a mutable quadruple.
It remains to assume that the five consecutive as above do not occur in $S$. It follows that $a$ can only occur among the first two elements or among the last two elements of $S$. Let $S^{\prime}$ be the sequence obtained by removal from $S$ the first two and the last two elements. Then each element of $S^{\prime}$ is contained in $I\setminus\{a\}$ and the the length of $S^{\prime}$ is at most $4(|I|-1)=4|I\setminus\{a\}|$. Therefore $S^{\prime}$, and hence $S$, contains a mutable quadruple by the induction assumption. ∎
In order to use Theorem 23, we map each colour interval of $P_{0}$ with a real number. The mapping is based on the invariance as in Lemma 25. Prior to that, we prove one more auxiliary lemma.
Lemma 24.
Let $U^{\prime}\subseteq Red\cup Blue$ be a context factor of $F$ and let $P^{\prime}$ be a colour interval of its spine. Then all the vertices of $P^{\prime}$ have the same foreign bags.
Proof.
Let $u_{1},u_{2}\in V(P^{\prime})$. Let $w\in fbag_{F}(u_{1})$. Since $u_{1}$ and $u_{2}$ are selected arbitrarily, it is enough to demonstrate that $w\in fbag_{F}(u_{2})$. As $w$ is not of the same colour as $u_{1}$, $w$ is a proper ancestor of $u_{1}$ and, in fact, $w$ is a proper ancestor of all the vertices of $P^{\prime}$. In particular, $w$ is a proper ancestor of $u_{2}$. Next, $w$ is not adjacent to $u_{1}$ since the vertices are of distinct colours. Therefore, $w$ is adjacent to a proper successor of $u_{1}$ and, in fact to a proper successor of all of $P^{\prime}$ (in light of Corollary 6). In particular, $u$ is adjacent to a proper successor of $u_{2}$. We conclude that $w\in bag_{F}(u_{2})$. Since $u_{2}$ and $w$ are of different colours, $w\in fbag_{F}(u_{2})$. ∎
With the notation as in Lemma 24, we introduce the notation $fbag_{F}(P^{\prime})$ as $fbag_{F}(u)$ for an arbitrary $u\in V(P^{\prime})$.
Lemma 25.
Let $P^{\prime}$ be a colour interval of $P_{0}$. Then all the vertices of $P^{\prime}$ have the same foreign balls.
Proof.
Let $u_{1},u_{2}\in V(P^{\prime})$. Let $W$ be a foreign component of $u_{1}$. Since vertices $u_{1}$ and $u_{2}$ are selected arbitrarily, it is enough to demonstrate that $W$ is a foreign component of $u_{2}$. It is immediate from Lemma 24 that $W\subseteq bag_{F}(u_{2})$. As $W$ is a connected component of $H[bag_{F}(u_{1})]$, $W$ is a connected set of $H[bag_{F}(u_{2})]$ but possibly not a maximal one. In this case, there is a vertex $w^{\prime}\in bag_{F}(u_{2})\setminus W$ such that $w^{\prime}$ is adjacent to $W$. We observe that $w^{\prime}$ cannot be the colour of $u_{2}$ since there $H$ has no hyperedges including vertices of two different colours. If $w^{\prime}$ is the colour of $W$ then by Lemma 24, $w^{\prime}\in bag_{F}(u_{1})$. But then $W\cup\{w^{\prime}\}$ is a connected subset of $H[bag_{F}(u_{1})]$ in contradiction to the maximality of $W$.
It remains to assume that $w^{\prime}\in X$. But then $W$ is a subset of the $X$-ball of $u_{2}$. But this is a contradiction since $W$ is disjoint with the $X$-ball of $u_{1}$ and the latter is equal to the $X$-ball of $u_{2}$ since both $u_{1}$ and $u_{2}$ belong to the same $X$-interval. ∎
We refer to $f\textsf{-width}(X-ball_{F}(u))$, $f\textsf{-width}(nball_{F}(u))$ and $f\textsf{-width}(fball_{F}(u))$ as the $X$-width, native width, and foreign width of $u$ respectively (w.r.t. $F$) and denote them by $X-width_{F}(u)$, $nwidth_{F}(u)$, and $fw_{F}(u)$  [^2] The naming $fwidth_{F}$, consistent with $nwidth_{F}$, could be easily confused with $f\textsf{-width}$. Therefore we chose another, easily distinguishable notation.. It follows from Lemma 25 that the foreign width is invariant for the vertices of a colour interval. Therefore, we use the notation $fw_{F}(P^{\prime})$ where $P^{\prime}$ is a colour interval to denote the foreign width w.r.t. $F$ of an arbitrary vertex of $P^{\prime}$.
Now, let $P^{\prime}_{1},\dots,P^{\prime}_{q}$ be the partition of $P_{0}$ into colour intervals that occur on $P_{0}$ in the order listed starting from the root of $U_{0}$. In the rest of the proof, we demonstrate that if a sequence $fw_{F}(P^{\prime}_{1}),\dots fw_{F}(P^{\prime}_{q})$ has a mutable quadruple then the desired neutral swap exists. We then harness Theorem 23 to demonstrate that if $q$ is sufficiently long w.r.t. $f\textsf{-width}(H)$ and $rank(H)$ then such a quadruple does exist.
Lemma 26.
Assume that there is $1\leq i\leq q-3$ so that
$fw_{F}(P^{\prime}_{i}),fw_{F}(P^{\prime}_{i+1}),fw_{F}(P^{\prime}_{i+2}),fw_{F}(P^{\prime}_{i+3})$ forms a mutable quadruple of $fw_{F}(P^{\prime}_{1}),\dots fw_{F}(P^{\prime}_{q})$. Then the swap of $P^{\prime}_{i+1}$ and $P^{\prime}_{i+2}$ is neutral.
Proof.
Throughout the proof, we will frequently refer to the set of predecessors and successors of a vertex in an elimination forest. We denote these sets by $pred_{F}(u)$ and $succ_{F}(u)$, respectively.
Let $F^{\prime}$ be the elimination forest resulting from the swap.
Let $u_{1}$ and $v_{1}$ be the respective first and last vertices of $P_{i+1}$, $u_{2}$ and $v_{2}$ be the respective first and last vertices of $P_{i+2}$ and $u_{3}$ be the first vertex of $P_{i+3}$. (the path are explored along the spine of $U_{0}$ starting from the root).
Claim 27.
Let $u\in(V(H)\setminus V(F_{u_{1}}))\cup V(F_{u_{3}})$. Then $bag_{F}(u)=bag_{F^{\prime}}(u)$
Proof of Claim. It is straightforward to observe that for each $u$ as in the statement of the claim, $pred_{F}(u)=pred_{F^{\prime}}(u)$ and $succ_{F}(u)=succ_{F^{\prime}}(u)$
Claim 28.
Let $u\in Cont_{F}(P_{i+1}+P_{i+2},\{u_{3}\})\setminus V(P_{i+1}\cup P_{i+2})$. Then $bag_{F}(u)=bag_{F^{\prime}}(u)$.
Proof of Claim. We start by observing that $succ_{F}(u)=succ_{F^{\prime}}(u)$. By Corollary 6, all the vertices of $succ_{F}(u)$ are of the same colour as $u$. Therefore, both $bag_{F}(u)$ and $bag_{F^{\prime}}(u)$ can have only elements of $X$ or elements of the same colour as $u$. But $pred_{F}(u)\cap(X\cup Native(u))=pred_{F^{\prime}}(u)\cap(X\cup Native(u))$.
In light of Claims 27 and 28, we only need to verify the lemma for $V(P_{i+1})\cup V(P_{i+2})$. Our goal is to demonstrate that for each $u\in V(P_{i+1})\cup V(P_{i+2})$, $X-ball_{F}(u)=X-ball_{F}(u)$, $nball_{F}(u)=nball_{F^{\prime}}(u)$ and $|fball_{F}(u)|\leq|fball_{F^{\prime}}(u)|$,
We observe (by construction and Corollary 6) that for each $u\in V(P_{i+1})\cup V(P_{i+2})$, $pred_{F}(u)\cap Native(u)=pred_{F^{\prime}}(u)\cap Native(u)$ and $succ_{F}(u)\cap(Native(u)\cup X)=succ_{F^{\prime}}(u)\cap(Native(u)\cup X)$.
We thus conclude that
Claim 29.
For each $u\in V(P_{i+1})\cup V(P_{i+2})$, $nbag_{F}(u)=nbag_{F^{\prime}}(u)$.
Claim 30.
$F^{\prime}$ is a reduced elimination forest. $U$ is a context factor of $F^{\prime}$. The sequence of colour intervals of $U$ in $F^{\prime}$ is the concatenation of ${\bf P}_{0}$, $F^{\prime}[V(P_{i})\cup V(P_{i+2})],F^{\prime}[V(P_{i+1})\cup V(P_{i+3}))]$, ${\bf P}_{1}$. Here ${\bf P}_{0}$ is a possibly empty sequence of colour intervals of $P$ occurring before $P_{i}$ and ${\bf P}_{1}$ is a possibly empty sequence of colour intervals of $P$ occurring after $P_{i+3}$.
Proof of Claim. The first two statements follow from Lemma 43. The third statement is immediate by construction. ∎
Claim 31.
For each $u\in V(P_{i+1})$, $fbag_{F^{\prime}}(u)=fbag_{F}(P_{i+3})$ and for each $u\in V(P_{i+2})$, $fbag_{F^{\prime}}(u)=fbag_{F}(P_{i})$.
Proof of Claim. We only prove the statement for $u\in V(P_{i+1})$, the other case is symmetric. Denote $F^{\prime}[V(P_{i+1})\cup V(P_{i+3})]$ by $P^{\prime}$. By Claim 30 and Lemma 24, $fbag_{F^{\prime}}(u)=fbag_{F^{\prime}}(P^{\prime})$. On the other hand $P^{\prime}$ includes vertices of $P_{i+3}$. Hence, by Claim 27, $fbag_{F^{\prime}}(P^{\prime})=fbag_{F}(P_{i+3})$. ∎
Claim 32.
For each $u\in V(P_{i+1})\cup V(P_{i+2})$, $X-bag_{F}(u)=X-bag_{F^{\prime}}(u)$.
Proof of Claim. We first observe that $pred_{F}(u)\cap X=pred_{F^{\prime}}(u)\cap X$. Let $u\in V(P_{i+2})$. By construction, $succ_{F^{\prime}}(u)=succ_{F}(u)\cup Cont_{F}(P_{i+1},\{u_{2}\})$. Therefore, if the claim is not true then $X-bag_{F^{\prime}}(u)$ is a strict superset of $X-bag_{F}(u)$. Let $v\in X-bag_{F^{\prime}}(u)\setminus X-bag_{F}(u)$. It follows that there is $w\in Cont_{F}(P_{i+1},\{u_{2}\})$ such that $v$ is adjacent to $w$. Let $u_{0}$ be an arbitrary vertex of $P_{i}$. We observe that $v\in pred_{F}(u_{0})$ and $w\in succ_{F}(u_{0})$. Hence, $v\in X-bag_{F}(u_{0})$. Due to $U_{0}$ being $X$-homogenous, $bag_{F}(u_{0})=bag_{F}(u)$ and hence $v\in bag_{F}(u)$ in contradiction to our assumption.
Now, let $u\in V(P_{i+1})$. Then $succ_{F^{\prime}}(u)=succ_{F}(u)\setminus Cont_{F}(P_{i+2},\{u_{3}\})$. Hence, if the claim does not hold, $X-bag_{F^{\prime}}(u)$ is a strict subset of $X-bag_{F}(u)$. Let $v\in X-bag_{F}(u)\setminus X-bag_{F^{\prime}}(u)$. Let $u_{0}\in P_{i+3}$. Due to the $X$-homogenicity of $U_{0}$, $v\in X-bag_{F}(u_{0})$. By Claim 27, $v\in bag_{F^{\prime}}(u_{0})$. Since $succ_{F^{\prime}}(u_{0})\subseteq succ_{F^{\prime}}(u)$, we conclude that $v\in X-bag_{F^{\prime}}(u)$, a contradiction. ∎
Claim 33.
For each $u\in V(P_{i+1})\cup V(P_{i+2})$, $X-ball_{F}(u)=X-ball_{F^{\prime}}(u)$.
Proof of Claim. In light of Claim 32, it is enough to observe that for each $x\in X-\mathsf{bag}_{F}(u)$, the connected component of $H[\mathsf{bag}_{F}(u)]$ containing $x$ is also a connected component of $\mathsf{bag}_{F^{\prime}}(u)$.
So, let $W$ be such a component. We need first to verify that $W\subseteq\mathsf{bag}_{F^{\prime}}$. $W\cap(nbag_{F}(u)\cup X-\mathsf{bag}_{F}(u))\subseteq\mathsf{bag}_{F^{\prime}}(u)$ by combination of Claims 29 and 32. Let $v\in W\cap fbag_{F}(u)$.
Assume first that $u\in V(P_{i+2})$. Let $u_{0}\in V(P_{i})$. Due to the $X$-homogenicity of $U_{0}$ and Claim 27, $v\in bag_{F^{\prime}}(u_{0})$. As $u$ and $u_{0}$ are of the same colour, $v\in fbag_{F^{\prime}}(u_{0})$. By Claim 31, $v\in fbag_{F^{\prime}}(u)\subseteq bag_{F^{\prime}}(u)$. If $u\in V(P_{i+1})$, we apply the same argument with $P_{i+3}$ being used instead of $P_{i}$.
By assumption about $W$, it is a connected set of $H$ and hence a connected set of $H[bag_{F^{\prime}}(u)]$ and hence, in turn, a subset of a connected component of $H[bag_{F^{\prime}}(u)]$. It remains to verify that there are no other vertices in this connected component. Such an extra vertex $v$ may only belong to $bag_{F^{\prime}}(u)\setminus bag_{F}(u)$. By combination of Claims 29 and 32, $v\in fbag_{F^{\prime}}(u)$. Assume that $u\in V(P_{i+2})$ Let $u_{0}\in V(P_{i})$. By Claim 31, $v\in bag_{F^{\prime}}(u_{0})$. On the other hand, by the $X$-homogenicity of $U_{0}$ and Claim 27, $W$ is a connected component of $H[bag_{F^{\prime}}(u_{0})]$, a contradiction. If $u\in V(P_{i+1})$ then the argument is symmetric with $P_{i+3}$ being used instead of $P_{0}$. ∎
Claim 34.
For each $u\in V(P_{i+1})$, $fball_{F^{\prime}}(u)=fball_{F}(P_{i+3})$ and for each $u\in V(P_{i+2})$, $fball_{F^{\prime}}(u)=fball_{F}(P_{i})$.
Proof of Claim. We prove the claim for $u\in V(P_{i+1})$, for the other case, the proof is symmetric. We first observe that $fball_{F^{\prime}}(u)=fbag_{F^{\prime}}(u)\setminus X-ball_{F^{\prime}}(u)$. Let $u_{0}\in V(P_{i+3})$ By Claim 31, $fbag_{F^{\prime}}(u)=fbag_{F^{\prime}}(u_{0})$. By combination of Claim 33, the $X$-homogenicity of $U_{0}$ and Claim 27, $X-ball_{F^{\prime}}(u)=X-ball_{F^{\prime}}(u_{0})$. We thus conclude that $fball_{F^{\prime}}(u)=fball_{F^{\prime}}(u_{0})$. By Claim 27, $fball_{F^{\prime}}(u)=fball_{F}(u_{0})$. Finally, by Lemma 25, $fball_{F^{\prime}}(u)=fball_{F}(P_{i+3})$. ∎
Claim 35.
For each $u\in V(P_{i+1})\cup V(P_{i+2})$, $nball_{F}(u)=nball_{F^{\prime}}(u)$.
Proof of Claim. We observe that $nball_{F^{\prime}}(u)=nbag_{F^{\prime}}(u)\setminus X-ball_{F^{\prime}}(u)=nbag_{F}(u)\setminus X-ball_{F}(u)=nball_{F}(u)$, the first inequality follows from the combination of Claims 31 and 33
Let $u\in V(P_{i+1})$. Then
$f(bag_{F^{\prime}}(u))$ $=$ $g_{f}(nwidth_{F^{\prime}}(u),fw_{F^{\prime}}(u),X-width_{F^{\prime}}(u))$
$=$ $g_{f}(nwidth_{F}(u),fw_{F}(P_{i+3}),X-width_{F}(u))$
$\leq$ $g_{f}(nwidth_{F}(u),fw_{F}(P_{i}+1),X-width_{F}(u))=f(bag_{F}(u),$
the first equality follows from Proposition 12, the second equality follows from the combination of Claims 35, 31, and 34, the inequality follows from the definition of a neutral swap.
For $u\in V(P_{i+2})$ the argument is the same with $P_{i}$ used instead of $P_{i+3}$ and $P_{i+2}$ used instead of $P_{i+1}$. ∎
Proof of Lemma˜16.
The following is immediate from the bounded number of values property.
Claim 36.
There is a function $\kappa_{2}$ such for hypergraph $H^{\prime}$ of rank at most $r$ and $U^{\prime}\subseteq V(H^{\prime})$ $f_{H^{\prime}}(U^{\prime})$ can take at most $\kappa_{2}(k,r)$ values smaller than or equal to $k$.
Let $Val$ be the set of all values $f_{H^{\prime}}(U^{\prime})$ that is at most $f\textsf{-width}(H)$, $U^{\prime}\subseteq V(H^{\prime})$, and $H^{\prime}$ is of rank at most $rank(H)$. According to Claim 36, $|Val|\leq\kappa_{2}(f\textsf{-width}(H),rank(H))$. Set $\kappa_{1}=4\kappa_{2}$ and assume that $q\geq\kappa_{1}(f\textsf{-width}(H),rank(H))$.
Clearly $fw_{F}(P^{\prime}_{1}),\dots fw_{F}(P^{\prime}_{q})$ is a sequence of elements of $Val$. By Theorem 23, there is $i\leq q-3$ such that $fw_{F}(P^{\prime}_{i}),fw_{F}(P^{\prime}_{i+1}),fw_{F}(P^{\prime}_{i+2}),fw_{F}(P^{\prime}_{i+3})$ is a mutable quadruple. By Lemma 26, the swap of $P^{\prime}_{i+1}$ and $P^{\prime}_{i+2}$ is neutral. We also note that none of these two is a terminal colour interval. ∎
A.3. Gathering the things together
The main idea is establishing a connection between the ‘red-blue’ case and the Reduced Dealteration lemma. First of all, we identify the ‘interface’ between two cases. In the red-blue case, we apply a series of swaps to obtain a reduced elimination forest with small monochromatic factorisations for both the red and the blue subsets. Apart from that, the resulting forest must preserve an invariant compared to the initial forest. To define this invariant, we introduce the notion of a replacement.
Definition 37.
A reduced elimination forest $F^{\prime}$ is a replacement of a reduced elimination forest $F$ (w.r.t. $(Red,X,Blue)$ if not clear from the context) if each monochromatic factor of $F$ is also a monochromatic factor of $F^{\prime}$. We say that $F^{\prime}$ is a strong replacement if $f\textsf{-width}(F^{\prime})\leq f\textsf{-width}(F)$.
From the reasoning perspective the statement serving as the interface between the red-blue case and the general case is what in (Bojańczyk and Pilipczuk, 2022) paper called Local Dealteration Lemma. We state the local version below.
Lemma 38.
There is a monotone function $\kappa_{2}$ such that the following holds. Let $F$ be a reduced elimination forest. Then there is a strong replacement $F^{\prime}$ of $F$ such that the maximal factorisation of both $Red$ and $Blue$ under $F^{\prime}$ is at most $\kappa_{2}(|X|,tw({\bf t}),f\textsf{-width}(F),rank(H),\Delta(H))$
We are going to prove Reduced Hypergraph Dealternation Lemma with Lemma 38 considered as given and then prove Lemma 38 itself. For the proof of the Reduced Hypergraph Dealternation Lemma we need three auxiliary lemmas that are stated below.
Lemma 39.
Let $W_{0}\subseteq W_{1}\subseteq V(H)$. Then $\textsc{mf}_{F}(W_{0})|\leq 9\cdot\textsc{mf}_{F}(W_{1})+3\cdot|W_{1}\setminus W_{0}|$.
Proof.
Lemma 7.10. of (Bojańczyk and Pilipczuk, 2022) proves this statement for graphs and it is not hard to upgrade the proof to hypergraphs. ∎
Lemma 40.
Let $F$ be a reduced elimination forest of $H$. Let $\ell$ be the treewidth of the tree decomposition indudced by $H$. Let $X,A_{1},\dots,A_{p}$ be a partition of $V(H)$ such that for $1\leq i\neq j\leq p$, no hyperedge of $H$ intersects both $A_{i}$ and $A_{j}$. Let $U\in\textsc{mf}(V(H)\setminus X)$ be a context factor. Then $U$ intersects at most $\ell+1$ sets among $A_{1},\dots A_{p}$.
Proof.
Lemma 7.5. of (Bojańczyk and Pilipczuk, 2022) proves this statement for graphs and it is not hard to upgrade the proof to hypergraphs. ∎
The following lemma is an upgrade of Lemma 7.11 of (Bojańczyk and Pilipczuk, 2022).
Lemma 41.
Let $({\bf t},F)$ be a tf-pair such that $F$ is a reduced elimination forest Then $mir({\bf t},F)$ is a function of $split({\bf t},F)$, $tw({\bf t}$), $f\textsf{-width}(F)$, and $rank(H)$.
Proof.
We demonstrate existence of a function $h$ so that ${\bf t}=(T,{\bf B})$ for an arbitrary node $x$ of $T$, the irregularity of $x$ is at most $h(split({\bf t},F),tw({\bf t}),f\textsf{-width}(F),rank(H))$.
Let $y_{1},\dots,y_{p}$ be the children of $x$. For each $1\leq i\leq p$, let $A_{i}=cmp(y_{i})$. Let $X=V(H)\setminus\bigcup_{i\in[p]}A_{i}$. Now, we employ Lemma 40 to observe that any maximal factor $B$ of $V(H)\setminus X$ that is a context factor intersects at most $\ell+1$ elements of $A_{1},\dots,A_{p}$ which is known to be a function of $f\textsf{-width}(F)$ and $rank(H)$.
We observe that the size of the maximal factorisation of $cmp(x)$ is at most $split({\bf t},F)$ simply by definition. We further note that $\bigcup_{i\in[p]}A_{i}=cmp(x)\setminus mrg(x)$. On the other hand, $mrg(x)\leq|bag(x)|$ which is at most $tw({\bf t})$. It follows from Lemma 39 that the maximal factorisation size of $A_{1}\cup\dots\cup A_{p}$ and hence the number of context factors in that factorisation, is upper bounded by a function of $split({\bf t},F)$ and $tw({\bf t})$. Let $I\subseteq[p]$ be the set of all $i$ such that $A_{i}$ intersects with a context factor of $\textsc{mf}(A_{1}\cup\dots\cup A_{p})$. Clearly $|I|$ is at most the upper bound of Lemma 40 multiplied by the number of context factors in $\textsc{mf}(A_{1}\cup\dots\cup A_{p})$. Hence $|I|$ is upper bounded by a function $h(split({\bf t},F),tw({\bf t}),f\textsf{-width}(F),rank(H))$.
We claim that if $A_{i}$ is not well-formed then $i\in I$ thus upper-bounding the irregularity of $x$ and hence the maximum irregularity since $x$ is taken arbitrarily. It remains to prove the claim. The union of tree and forest factor in the factorisation of $\bigcup_{i\in[p]}A_{i}$ can be seen as the union of tree factors $B_{1},\dots,B_{r}$. Let $i\in[p]\setminus I$. We are going to prove that $A_{i}$ is regular which is the same as to say that the factorisation of $A_{i}$ consists of the union of tree and forest factors.
Let $J\subseteq[r]$ be the set such that $A_{i}\subseteq\bigcup_{j\in J}B_{j}$. We observe that for each $j\in J$, $B_{j}\subseteq A_{i}$, Indeed, $B_{j}\subseteq A_{1}\cup\dots\cup A_{p}$. On the other hand, since $H[B_{j}]$ is connected, $B_{j}$ cannot belong to several distinct $A_{i}$. It follows that $A_{i}=\bigcup_{j\in J}B_{j}$. Now, assume for the sake of contradiction that a maximal factor $C$ of $A_{i}$ is context one. Let $u$ be the root of $C$. As we have observed $u\in B_{j}$ for some $j\in J$. But then $F_{u}\subseteq B_{j}$. That is $C\subset F_{u}\subseteq B_{j}$ in contradiction to $C$ being maximal. ∎
Proof of the Reduced Hypergraph Dealternation Lemma (Lemma 7).
First, we assume that $F$ is reduced because the transformation from an arbitrary forest to a reduced one can be done so that each bag of the tree decomposition induced by the resulting forest is a subset of a bag of the tree decomposition induced by the original forest (Lemma 7.8 of (Bojańczyk and Pilipczuk, 2022)). Consequently, the $f\textsf{-width}$ of the resulting reduced forest does not increase.
The next step is to define canonical Red-Blue decompositions w.r.t. the nodes $x$ of $T$ (recall that ${\bf t}=(T,B)$). In particular, the decomposition w.r.t $x$ is $(cmp(x),adh(x),V(H)\setminus(cmp(x)\cup adh(x)))$. We let $(x_{1},\dots,x_{m})$ be the sequence of nodes of $T$ being explored in a bottom up manner, that is children occur in this sequence before their parents. We observe the following.
Claim 42.
Let $1\leq i<j\leq m$. Then $cmp(x_{i})$ is monochromatic for the canonical red-blue decomposition w.r.t. $x_{j}$.
Proof of Claim.
If $x_{j}$ is a predecessor of $x_{i}$ then $cmp(x_{i})\subseteq cmp(x_{j})$ hence the claim holds. Otherwise, we observe that $cmp(x_{i})\subseteq V(H)\setminus(adh(x_{j})\cup cmp(x_{j}))$. Indeed $cmp(x_{i})$ does not intersect with $bag(x)$ if $x$ is not a successor of of $x_{i}$. We note that $adh(x_{j})\cup cmp(x_{j})$ is a subset of the union of $bag(x_{j})$ and the bags of the proper successors of $x_{j}$. By assumption $x_{j}$ is not a predecessor of $x_{i}$. By assumption about $(x_{1},\dots,x_{m})$ $x_{j}$ is not a successor of $x_{i}$ either. It follows that no successor of $x_{j}$ is a successor of $x_{i}$. Hence the union of the bags of successors of $x_{j}$ is disjoint with $cmp(x_{i})$ as required. ∎
Let $F_{0}=F$ and let $F_{1},\dots\ F_{m}$ be reduced elimination forests of $H$ defined as follows. For each $1\leq i\leq m$, $F_{i}$ is a strong replacement of $F_{i-1}$ with respect to the canonical red-blue decomposition for $x_{i}$. We are going to prove by induction on $1\leq i\leq m$ that, for each $j\leq i$, $comp(x_{j})$ has maximal factorisation of size at most $\kappa_{3}(tw({\bf t}),f\textsf{-width}(F),rank(H),\Delta(H))=\kappa_{2}(|X|,tw({\bf t}),f\textsf{-width}(F),rank(H),\Delta(H))$.
Let $i=1$. Then this holds for $comp(x_{1})$ simply by Lemma 38. Since $i$ is the smallest index, the remaining part of the statement is vacuously true. Assume first that $i>1$. First, let $j=i$. Then the statement for $j$ is immediate from Lemma 38. Assume now that $j<i$. By the induction assumption, the statement about $x_{j}$ is true regarding $F_{i-1}$. As $F_{i}$ is a replacement of $F_{i-1}$ and, by Claim 42, $comp(x_{j})$ is monochromatic for the canonical red-blue decomposition for $x_{i}$, a witnessing factorisation for $comp(x_{j})$ w.r.t. $F_{i-1}$ remains so w.r.t. $F_{i}$.
As $|X|$ in the considered canonical decompositions is at most the treewidth of the tree decomposition induced by $F$, which is, in turn upper bounded by a function of $f\textsf{-width}(F)$ and $rank(H)$, it follows from the above reasoning that
$split({\bf t},F_{m})\leq\kappa_{3}(tw({\bf t}),f\textsf{-width}(F),rank(H),\Delta(H))$
as required. It remains to add that $mir({\bf t},F_{m})$ is also upper-bounded by a function of the same parameters by Lemma 41. ∎
It now remains to prove Lemma 38. First, we need an auxiliary lemma specifying properties of a single swap.
Lemma 43.
Let $F$ be a reduced elimination forest. Let $B$ be a context factor of $F$ with the spine $P$ and appendices $W$. Let $P_{1},\dots,P_{q}$ be the colour intervals of $P$. Let $1\leq i\leq q-4$. Let $F^{\prime}$ be the elimination forest obtained from $F$ by swapping $P_{i+1}$ and $P_{i+2}$. Then the following statements hold regarding $F^{\prime}$.
  1. (1)
    $F^{\prime}$ is a reduced elimination forest.
  2. (2)
    Each monochromatic factor of $F$ remains a monochromatic factor of $F^{\prime}$ in the same capacity (that is a tree factor remains a tree factor, a forest factor remains a forest factor, and a context factor remains a context factor).
  3. (3)
    Each context factor of $F$ disjoint with $B$ remains a context factor in $F^{\prime}$ with the same spine and appendices.
  4. (4)
    $B$ remains a context factor of $F^{\prime}$ with the same appendices but with the spine obtained from $P$ by swapping $P_{i+1}$ and $P_{i+2}$ as in Definition 7. In particular, the number of colour intervals of $B$ w.r.t. $B$ reduces by $2$.
Proof.
First, it is not hard to observe that
Claim 44.
$F$ and $F^{\prime}$ have the same set of leaves.
Suppose that $F^{\prime}$ is not a reduced elimination forest. This means that $F^{\prime}$ has an internal vertex $u$ such that $u$ is not adjacent to $F^{\prime}_{u}\setminus\{u\}$. By Claim 44, $u$ is an internal vertex of $F$. Let $v$ be a neighbor of $u$ in $F_{u}$. Then $v\in Native(u)\cup X$. By construction, $v\in F^{\prime}_{u}$ and hence $u$ is adjacent to $F^{\prime}_{u}\setminus\{u\}$, a contradiction.
We further observe that if $F_{u}$ is monochromatic then for each vertex of $F_{u}$, its adjacency is invariant in $F^{\prime}$. Hence the monochromatic tree factors of $F$ remain ones in $F^{\prime}$. Further on, we notice that if $F_{u}$ is monochromatic then its parent remains the same in $F^{\prime}$. We thus conclude that monochromatic tree factors of $F$ remain ones in $F^{\prime}$.
In order to demonstrate that monochromatic context factors are invariant in $F^{\prime}$, we first observe the following.
Claim 45.
A monochromatic path of $F$ remains a monochromatic path of $F^{\prime}$.
Now, let $U$ be a monochromatic context factor of $F$ and let $Q$ be its spine. By Claim 45, $Q$ is monochromatic in $F^{\prime}$. Further on, again by all Claim 45, for each vertex $v$ of $Q$, the intersection of the neighborhood of $v$ with $Native(v)$ remains invariant in $F^{\prime}$. Taking into account the invariance of the monochromatic tree factors rooted by the children of the vertices of $Q$, we conclude that $F[U]=F^{\prime}[U]$. Finally, the final vertex if $Q$ is not a leaf by Claim 44. We conclude that $U$ remains a monochromatic context factor in $F^{\prime}$.
Let $U$ be a context factor of $F$ disjoint with $B$. We observe that, by construction $F[U]=F^{\prime}[U]$ and the neighborhood of the spine of $U$ also invariant. Hence, the penultimate statement of the lemma holds.
The last item of the lemma is immediate by construction. ∎
Definition 46.
Let $F$ be a reduced elimination forest of $H$ and let ${\bf U}=\{U_{1},\dots U_{q}\}$ be a factorisation of $Red\cup Blue$ w.r.t. $F$ where each $U_{i}$ is a monochromatic tree or forest factor or a (not necessarily monochromatic) context factor. A reduced elimination forest $F^{\prime}$ is an strong replacement of $F$ w.r.t. $\bf U$} if the following conditions hold.
  1. (1)
    $f\textsf{-width}(F^{\prime})\leq f\textsf{-width}(F)$.
  2. (2)
    Each monochromatic factor of $F$ remains a monochromatic factor in the same capacity (that is a tree factor remains a tree faxctor, a forest factor remains a forest factor, and a context factor remains acontext factor).
  3. (3)
    Each context factor of $F$ that is an element of $\bf U$} remains a context factor in $F^{\prime}$.
Remark 47.
Let us make three remarks related to the notion of a strong replacement. First, if $F^{\prime}$ is a strong replacement of $F$ w.r.t. to a factorisation $\bf U$} then $\bf U$} is also a factorisation of $F^{\prime}$ where each factor retains its capacity. Second, the relation is transitive. That is, if $F^{\prime\prime}$ is a strong replacement of $F^{\prime}$ w.r.t. $\bf U$} then $F^{\prime\prime}$ is a strong replacement of $F$ w.r.t. $\bf U$}. Finally, $F^{\prime}$ is a strong replacement of $F$ in the sense of Definition 37.
Lemma 48.
Let $F$ be a reduced elimination forest of $H$ and let ${\bf U}=\{U_{1},\dots U_{q}\}$ be a factorisation of $Red\cup Blue$ w.r.t. $F$ where each $U_{i}$ is a monochromatic tree or forest factor or a (not necessarily monochromatic) context factor. Then there is strong replacement $F^{\prime}$ of $H$ where the number of colour intervals in the spine of each context factor $U\in{\bf U}$ is at most $\kappa(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$ where $\kappa$ is the function as in Lemma 8.
Proof.
For each $i\in[q]$ define the excess $x_{i}$ as follows. If $U_{i}$ is not a context factor or if $U_{i}$ is a context factor but the number $y_{i}$ of colour intervals of its spine is at most $\kappa(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$ then $x_{i}=0$. Otherwise, we let $x_{i}=y_{i}-\kappa(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$. We prove the lemma by induction on $x=sum_{i\in[q]}x_{i}$ which we refer to as the excess of $\bf U$} w.r.t. $F$. If $x=0$ then simply set $F^{\prime}=F$.
So, we assume that $x>0$. We are going to demonstrate existence of a strong replacement $F^{*}$ of $F$ w.r.t. $\bf U$} such that the excess of $\bf U$} w.r.t. $F^{*}$ is smaller than w.r.t. $F$. The lemma will then follow from the induction assumption combined with the transitivity of strong replacement w.r.t. $\bf U$} as specified in Remark 47.
Since $x>0$, we identify a context factor $U\in{\bf U}$ with the number of colour intervals greater than $\kappa(|X|,f\textsf{-width}(H),rank(H),\Delta(H))$. By Lemma 8, there is a neutral an elimination forest $F^{*}$ obtained from $F$ by a neutral swap of colour intervals of $U$. We first note that $f\textsf{-width}(F^{*})\leq f\textsf{-width}(F)$ simply by definition of a neutral swap. Then we employ Lemma 43 to conclude that $F^{*}$ is a reduced elimination forest and that the remaining properties of a strong replacement w.r.t. $\bf U$} hold as well. It also follows from Lemma 43 that the number of colour intervals of $U$ w.r.t. $F^{*}$ reduced by $2$ and the number of such intervals in the other context factors of $\bf U$} remains the same. Thus we conclude that the excess of $\bf U$} w.r.t. $F^{*}$ is smaller than w.r.t. $F$ as required. ∎
Proof of Lemma˜38.
[…]
Appendix B Verifying Hypergraph Widths in MSO
Let $\mathbf{H}_{f,k,r}$ be the set of all hypergraphs $H$ of rank at most $r$ such that $f_{H}(V(H))\leq k$. Let $\mathit{Val}_{f,k,r}=\{f_{H}(V(H))\mid H\in{\bf H}_{f,k,r}\}$.
Lemma 1.
There is an algorithm with runtime upper bounded by a monotone function on $k$ and $r$ that returns $\mathbf{H}_{f,k,r}$ and $\mathit{Val}_{f,k,r}$.
Proof.
Let $\beta=\beta[f]$. By the bounded size property, $\mathbf{H}_{f,k,r}$ is a subset of the set of hypergraphs of rank at most $r$ and with at most $\beta(k,r)$ vertices. Moreover, by the automatizability property, the value $f_{H}(V(H))$ for each such a graph can be tested by an algorithm upper bounded by a monotone function of $k$ and $r$. Therefore $\mathbf{H}_{f,k,r}$ and $\mathit{Val}_{f,k,r}$ can be generated by the following simple algorithm. Start from two empty sets (one for storage of the elements of $\mathbf{H}_{f,k,r}$, the other for $\mathit{Val}_{f,k,r}$). Generate all hypergraphs $H$ of rank at most $r$ and with at most $\beta(k,r)$ vertices. For each such a graph $H$, test $f_{H}(V(H))$. If $f_{H}(V(H))$ is at most $k$, add $H$ to the set for graphs and add $f_{H}(V(H))$ to the set for values. At the end of the exploration, return the resulting sets. ∎
Lemma 2.
Let $H$ be a hypergraph of rank at most $r$ and let $U\subseteq V(H)$. Then $f_{H}(U)\leq k$ if and only if $H[U]\in{\bf H}_{f,k,r}$.
Proof.
We simply observe that, by the invariant locality property, $f_{H}(U)=f_{H[U]}(V(H[U]))$. Then, simply by definition of ${H}_{f,k,r}$, $f_{H[U]}(V(H[U])\leq k$ if and only if $H[U]\in{\bf H}_{f,k,r}$. ∎
Lemma 3.
Let $H$ be a hypergraph with $q$ vertices and $r$ edges. Let $M$ be a model over a signature containing relations $Vertex,Edge,Adjacent$ where $Vertex$ and $Edge$ are unary relations and $Adjacent$ is a binary relation. such that $Vertex(M)$ and $Edge(M)$ are disjoint and $Adjacent(M)$ consists of elements $(v_{i},e_{j})$ with $v_{i}\in Vertex(M),e_{j}\in Edge(M)$. In other words, $M$ models a hypergraph $H(M)$ with $(Vertex(M),Edge(M),Adjacent(M))$ being the incidence graph with the roles of the sets obvious from their names.
Then there is an MSO formula $\varphi_{H}(U)$ such that $M\models\varphi_{H}(U)$ if and only if $U\subseteq Vertex$ and $H(M)[U]$ is isomorphic to $H$.
Proof.
First we arbitrarily enumerate the vertices of $H$ as $v_{1},\cdots,v_{q}$ and edges of $H$ as $e_{1},\cdots,e_{r}$. We let $Inc(H)$ be the incidence relation of $H$ consisting of pairs $(i,j)$ such that vertex $i$ belongs to edge $j$.
Now define the formula
$\varphi_{H}(U):=\ Vertices(U)\land\exists v_{1},\cdots,v_{q},e_{1},\cdots,e_{r}\ ($ $IsSet(U,v_{1},\cdots,v_{r})$
$\land AreEdges(U,e_{1},\cdots,e_{r})$
$\land Incidence_{H}(v_{1},\cdots,v_{q},e_{1},\cdots,e_{r}))$
with the conjuncts defined as follows. $Vertices(U)$ is true if and only if $U\subseteq Vertex$, implemented as $Vertices(U):=\forall u:U(u)\to\mathit{Vertex}(u)$.
The formula $IsSet(U,v_{1},\cdots,v_{q})$ is true if and only if $U=\{v_{1},\cdots,v_{q}\}$. We implement this by testing the following three properties.
  • $AllIn(U,v_{1},\cdots,v_{q})$ checks that each $v_{i}$ belongs to $U$. In terms of MSO, $AllIn(U,v_{1},\cdots,v_{q}):=U(v_{1})\land\cdots\land U(v_{q})$.
  • $NoneElse(U,v_{1},\cdots,v_{q})$ checks that each element of $U$ is one of $v_{1},\cdots,v_{q}$. In terms of MSO, $NoneElse(U,v_{1},\cdots,v_{q}):=\forall u:U(u)\to(u=v_{1}\lor\cdots\lor u=v_{q})$.
  • $AllDiff(v_{1},\cdots,v_{q})$ checks that all $v_{i}$ are distinct. In terms of MSO, $AllDiff(v_{1},\cdots,v_{q}):=\bigwedge_{1\leq i<j\leq q}v_{i}\neq v_{j}$.
It is not hard to see that $IsSet(U,v_{1},\cdots,v_{q}):=AllIn(U,v_{1},\cdots,v_{q})\land NoneElse(U,v_{1},\cdots,v_{q})\land AllDiff(v_{1},\cdots,v_{q})$ has the required properties. The formula $AreEdges(U,e_{1},\cdots,e_{r})$ is true if and only if the subset of edges adjacent to at least one element of $U$ in the adjacency relation is precisely $\{e_{1},\cdots,e_{r}\}$. To put it differently, $\{e_{1},\cdots,e_{r}\}$ are the only hyperedges of $H$ having non-empty intersection with $U$. We implement the formula as the conjunction of the following components.
  1. (1)
    $AllIntersect(U)$, which is true if and only if each of $e_{1},\cdots,e_{r}$ intersects $U$ (being considered as a hyperedge of $H(M)$). For this, we first need an auxiliary function $Intersect(U,e)$ (which will also be useful later on) testing whether a specific hyperedge intersects $U$). In terms of MSO, $Intersect(U,e):=\exists u:U(u)\land Adjacent(u,e)$. Now, $AllIntersect(U,e_{1},\cdots,e_{r}):=\bigwedge_{i\in[r]}Intersect(U,e_{i})$.
  2. (2)
    $NonElseInter(U,e_{1},\cdots,e_{r})$, which is true if and only if no hyperedge but $e_{1},\cdots,e_{r}$ intersects $U$. In terms of MSO, $NoneElseInter(U,e_{1},\cdots,e_{r}):=\forall e:Intersect(U,e)\rightarrow\bigvee_{i\in[r]}e=e_{i}$.
  3. (3)
    That $e_{1},\cdots,e_{r}$ are all distinct is tested by the already defined $AllDiff(e_{1},\cdots,e_{r})$.
It is not hard to see that
$AreEdges(U,e_{1},\cdots,e_{r}):=\$$AllIntersect(U,e_{1},\cdots,e_{r})$
$\land NoneElseInter(U,e_{1},\cdots,e_{r})$
$\land AllDiff(e_{1},\cdots,e_{r})$
has the required properties.
The formula $Incidence_{H}(v_{1},\cdots,v_{q},e_{1},\cdots,e_{r})$ is true if and only if the adjacency relation between $v_{1},\cdots,v_{q}$ and $e_{1},\cdots,e_{r}$ in $H(M)[U]$ is precisely as the adjacency relation between $[q]$ and $[r]$ in $H$ under the natural mapping. In particular, $Incidence_{H}(v_{1},\cdots,v_{q},e_{1},\cdots,e_{r}):=\bigwedge_{(i,j)\in Inc(H)}Adjacent(v_{i},e_{j})\land\bigwedge_{i\in[q],j\in[r],(i,j)\notin Inc(H)}\neg Adjacent(v_{i},e_{j})$. The first conjunct checks that all the incidence relations as in $H$ are present in $H(M)[U]$ and the second conjunct checks that no other incidence pairs are present are in $H(M)[U]$. Note a subtlety, for the second conjunct we need to restrict quantification over the indices of vertices to be within $[q]$ and for the indices of hyperedges to be within $[r]$.
Assume now that $H(M)[U]$ is isomorphic to $H$. This means that it is possible to enumerate the vertices of $U$ as $v_{1},\cdots,v_{q}$ and the edges of $H(M)[U]$ as $e_{1},\cdots,e_{r}$ so that $v_{i}$ is contained in $e_{j}$ in $H(M)[U]$ if and only if $i$ is contained in $j$ in $H$. Clearly, this means that $\varphi_{H}(U)$ is true. Conversely, suppose that $\varphi_{H}(U)$ is true. Then there is an enumeration $v_{1},\cdots,v_{q}$ of the vertices of $U$ and enumeration $e_{1},\cdots,e_{r}$ of the hyperedges of $H_{M}$ that have the same adjacency as $[q]$ and $[r]$ for $H$ under the natural mapping. This means that $H(M)[U]$ is isomorphic to $H$ under the mappings $v_{i}\to i$ for $i\in[q]$ and $e_{j}\to j$ for $j\in[r]$. ∎
Lemma 4.
Let $\mathbf{H}$ be a set of hypergraphs such that there is a $q$ such that for each $H\in\mathbf{H},|V(H)|\leq q$. Then there exists an MSO formula $\varphi_{\mathbf{H}}(U)$ such that $M\models\varphi_{\mathbf{H}}(U)$ if and only if $U\subseteq Vertices(M)$ and $H(M)[U]$ is isomorphic to an element of $\mathbf{H}$.
Proof.
As the number of non-isomorphic graphs of at most $q$ vertices is finite, $|\mathbf{H}|$ is finite. Therefore we utilize Lemma 3 and set $\varphi_{\mathbf{H}}(U):=\bigvee_{H\in\mathbf{H}}\varphi_{H}(U)$. ∎
Proof of Lemma˜11.
Let ${\bf H}={\bf H}_{f,k,r}$ It follows from the combination of Lemma 2 and Lemma 4 that $\varphi_{\bf H}(U)$ (as in Lemma 4) is true if and only if $f_{H(M)}(U)\leq k$.
Now, $\varphi^{*}_{f,k,r}:=\forall u:Vertex(u)\rightarrow\exists U:IsBag(u,U)\land\varphi_{\bf H}(U)$, where $M\models IsBag(u,U)$ if and only if $U=bag_{F(M)}(u)$. In terms of MSO, the formula can be defined as follows. $IsBag(u,U):=\forall v:U(v)\leftrightarrow u=v\lor Vertex(v)\land Descendant(v,u)\land\exists wDescendant(u,w)\land Neighb(v,w)$ where $Descendant(u_{1},u_{2})$ is true if and only there is a path from $u_{1}$ to $u_{2}$ in $(Vertices,Child)$ (well known to be MSO definable) and $M\models Neighb(v,w)$ if and only $v$ and $w$ are adjacent in $H(M)$. In terms of MSO, this can be defined as follows. $Neighb(v,w):=\exists e:Edge(e)\land Adjacent(v,e)\land Adjacent(w,e)$.
The second statement of the lemma follows from the combination of Lemma 1, and the construction of the MSO formulas as described in the proofs of lemma 3, 4, and the above part of the present proof.
The last statement of the lemma follows from the combination of Lemma 1 and 2. ∎
Appendix C Proof of Theorem 10
Definition 1.
With the notation as in Definition 8 the stain intersection graph $\mathsf{SG}({\bf t},F)$ is a bipartite graph with parts corresponding to $V(T)$ and $V(H)$ with $x\in V(T)$ adjacent to $u\in V(H)$ if and only if $x\in Stain(u,({\bf t},F))$.
Theorem 2.
There is a function $h$ such that the max-degree of a vertex of $V(T)$ in $\mathsf{SG}({\bf t},F)$ is at most $h(split({\bf t},F),mir({\bf t},F),f\textsf{-width}({\bf t}),\mathit{rank}(H))$.
Proving Theorem 2 requires some preparation.
Definition 3.
Let $x\in V(T)$ and let $\{x,u\}\in E(\mathsf{SG}({\bf t},F))$. We say that $\{x,u\}$ is a margin edge if $u\in mrg(x)$.
Definition 4.
Let $x\in V(T)$. The characteristic partition of $x$ denoted by $\textsc{cp}_{\bf t}(x)$ (the subscript may be omitted if clear from the context) if the family of all non-empty sets among
$\{mrg(x),V(H)\setminus cmp(x)\}\cup\{cmp(y)|y\in children_{T}(x)\}.$
Further on, the characteristic factorisation of $x$ denoted by $\textsc{cf}_{{\bf t},F}(x)$ (again, the subscript may be omitted if clear from the context) is $\bigcup_{U\in\textsc{cp}(x)}\textsc{mf}_{F}(U)$.
Lemma 5.
The number of context factors of $\textsc{cf}_{{\bf t},F}(x)$, that are not subsets of $mrg(x)$, is at most $mir({\bf t},F)\cdot split({\bf t},F)+3\cdot split({\bf t},F)$.
Proof.
Each context factor $U$ as specified in the statement can be unambiguously mapped to an element $U^{*}$ of $\textsc{cp}_{\bf t}(x)$ such that $U\subseteq U^{*}$. We observe that the number of such $U^{*}$ is at most $mir({\bf t},F)+1$. Indeed, $U^{*}$ can be $cmp(y)$ such that $y$ is a child of $x$ (and there are at most $mir({\bf t},F)$ such children) or $U^{*}$ may be $V(H)\setminus cmp(x)$. Each element $cmp(y)$ has at most $split({\bf t},F)$ factors in its maximal factorisation, so the total number of context factors yielded as subsets of $cmp(y)$ for $y$ being a child of $x$ is $mir({\bf t},F)\cdot split({\bf t},F)$. Further on, the total number of factors in the maximal factorisation of $cmp(x)$ is $split({\bf t},F)$, so the number of factors in the maximal factorisation of its complement is at most $3\cdot split({\bf t},F)$ by Lemma 3.3. of (Bojańczyk and Pilipczuk, 2022). Clearly, the number of context factors in the factorisation is at most this amount. ∎
Lemma 6.
Let $x\in V(T)$, $u\in V(G)$ and $\{x,u\}\in\mathsf{SG}({\bf t},F)$. Suppose that $\mathit{mn}(u)\neq x$ (or, to put it differently, $u\notin mrg(x)$). Let $U\in\textsc{cf}_{{\bf t},F}(x)$ be the element containing $u$. Then $U$ is a context factor of $F$ where $u$ is the parent of the appendices.
Proof.
By assumption, the only way $x$ to be included in $Stain(u,({\bf t},F))$ is to be not a first vertex on a path between $\mathit{mn}(u)$ and $\mathit{mn}(v)$ where $v$ is a child of $u$ in $F$. Let $W_{0}$ and $W_{1}$ be elements of $\textsc{cp}_{{\bf t},F}(x)$ containing $u$ and $v$ respectively.
Claim 7.
$W_{0}\neq W_{1}$.
Proof.
If $W_{1}=mrg(x)$, we are done by assumption. Next, assume that $W_{1}=cmp(y)$ where $y$ is a child of $x$. It follows then that $\mathit{mn}(v)\in V(T_{y})$. If we assume that $W_{0}=W_{1}$ then $\mathit{mn}(u)\in V(T_{y})$. Hence, the path between $\mathit{mn}(u)$ and $\mathit{mn}(v)$ fully lies in $T_{y}$ and does not involve $x$, a contradiction. It remains to assume that $W_{1}=V(G)\setminus cmp(x)$. Again if we assume that $W_{0}=W_{1}$ then both $\mathit{mn}(u)$ and $\mathit{mn}(v)$ lie outside of $T_{x}$ and hence the path between them does not involve $x$ either causing the same contradiction. ∎
As $U\subseteq W_{0}$, it follows from the claim that $u$ is contained in a factor while its child is not. It can only be possible if $U$ is a context factor and $u$ is the appendices parent. ∎
Proof of Theorem˜2.
The number of margin edges incident to $x$ is at most the margin size of $x$ which is in turn at most the bag size of $\bf t$} which, again in turn, is upper bounded by a function of $f\textsf{-width}({\bf t})$ and $\mathit{rank}(H)$.
Let $NM(x)$ be the neighbours of $x$ in $\mathsf{SG}({\bf t},F)$ connected to $x$ through a non-margin edge. Let ${\bf y}:NM(x)\rightarrow\textsc{cf}_{{\bf t},F}(x)$ be mapping each vertex of $NM(x)$ to the factor containing it. By Lemma 6, each $u\in NM(x)$ is mapped to a context factor where $u$ is the appendices’ parent. As each context factor contains precisely one appendices’ parent this mapping is injective. It follows that $NM(x)$ is upper bounded by the number of context factors in $\textsc{cf}_{{\bf t},F}(x)$ whse size is, in turn, upper bounded by a function of $split({\bf t},F)$ and $mir({\bf t},F)$ by Lemma 5.
Proof of Theorem˜10.
The proof consists of the following steps.
  1. (1)
    The graph $\mathsf{CG}({\bf t},F)$ is chordal as it is the intersection graph of connected subgraphs of a forest (Golumbic, 2004).
  2. (2)
    It follows from the chordality of $\mathsf{CG}({\bf t},F)$ that its chromatic number is at most the size of its maximal clique (Golumbic, 2004).
  3. (3)
    It is known that the intersection graph of connected subgraphs of a forest has the Helly property. In other words, if all pairwise intersection of stains are non-empty, then there is a vertex that belongs to all the stains.
  4. (4)
    It follows that the maximum clique size of $\mathsf{CG}({\bf t},F)$ is upper bounded by the max-degree of a vertex of $V(T)$ in $\mathsf{SG}({\bf t},F)$. Indeed, let $u_{1},\dots u_{q}$ form a clique of $\mathsf{CG}({\bf t},F)$. This means that for each $i\neq j\in[q]$, $Stain(u_{i},({\bf t},F))\cap Stain(u_{j},({\bf t},F))\neq\emptyset$. It follows from the Helly property that $\bigcap_{i\in[q]}Stain(u_{i},({\bf t},F))\neq\emptyset$. Let $x\in\bigcap_{i\in[q]}Stain(u_{i},({\bf t},F))$. Then $x$ is adjacent to each of $u_{1},\dots,u_{q}$ in $\mathsf{SG}({\bf t},F)$.
  5. (5)
    The theorem is now immediate from Theorem 2.
Appendix D Proof of Lemma 9
Following (Bojańczyk and Pilipczuk, 2022), we define an elementary MSO transduction as
one of the following relations.
Colouring.:
Colouring adds a unary predicate to the input structure, i.e. the output vocabulary has an additional unary predicate.
The input model $M$ is mapped to the set of all possible models obtained from $M$ by adding a unary relation over the universe of $M$. We note that colouring is the only guessing transaction mapping the input model to several output models.
MSO Interpretation.:
For any relation name $R$ in the output, there exists a formula $\varphi_{R}(x_{1},\cdots,x_{k})$ with the same arity as $R$. For a model $M$, the output is a singleton whose only element has the same universe as $M$ in addition to relations $R$ such that $(x_{1},\cdots,x_{k})\in R$ if and only if $\varphi_{R}(x_{1},\cdots,x_{k})$ is true. We note that the MSO interpretation is the elementary transduction that allows an arbitrary signature change (the colouring also changes the signature but only in a specific way of adding a single unary relation).
Copying.:
We do not use this elementary transduction for the proof below, so we do not provide further technical elaboration.
Filtering.:
For an MSO formula $\varphi$ there is a filtering transduction which filters out structures which do not satisfy $\varphi$, i.e. a structure $M$ is mapped to $\{M\}$ if it satisfies $\varphi$ and to $\{\}$ otherwise. This is the elementary transduction that can cause a model to be mapped to an empty set.
Universe restriction.:
For a given formula $\varphi(x)$ with one free first-order variable, the input $M$ is mapped to a singleton containing a structure whose universe is restricted to those elments $x$ that satisfy $\varphi$
The central part of the proof is the notion of a transduction tuple as defined below.
Definition 1.
Let $({\bf t}=(T,B),F)$ be a tf pair of a hypergraph $H$. A transduction tuple $({\bf C},{\bf D},{\bf K},{\bf f},{\bf h})$ is defined as follows.
  1. (1)
    $\bf C$} is a partition of $V(H)$ into independent sets of $\mathsf{CG}({\bf t},F)$. To put it differently $\bf C$} determines a proper colouring of $\mathsf{CG}({\bf t},F)$.
  2. (2)
    $\bf D$} is a weak partition of $V(H)$.
  3. (3)
    $\bf K$} is a family of subgraphs.
  4. (4)
    $\bf f$} is a mapping from $\bf C$} to $\bf D$} defined as follows.
    1. (a)
      There is exactly one element $D_{0}\in{\bf D}$ such that ${\bf f}^{-1}(D_{0})=\emptyset$.
    2. (b)
      For each $C\in{\bf C}$, ${\bf f}(C)$ is the set of all children (in $F$) of the non-leaf elements of $C$.
  5. (5)
    $\bf g$} is a surjective mapping from $\bf C$} such that for each $C\in{\bf C}$, ${\bf g}(C)=\bigcup_{u\in C}T[\mathit{Stain}(u,({\bf t},F))]$.
We also introduce additional related terminology. In particular, for each $u\in V(H)$, $C(u)$ is the element of $\bf C$} containing $u$. Also, for each $C\in{\bf C}$, we define a function ${\bf h}_{C}$ on $C$ such that, for each $u\in C$, ${\bf h}_{C}(u)$ is the connected component of ${\bf g}(C)$ containing $u$.
In the subsequent proof, we demonstrate that, on the one hand, transduction tuple, together with $\bf t$} allows reconstruction of $F$ and, on the other hand, it can be guessed by a transduction. For the reconstruction of $F$, we need the following definition.
Definition 2.
With the notation as in Definition 1, let childchild be a binary relation on $V(H)$ consisting of all pairs $(u,v)$ such that the following is true.
  1. (1)
    $v\in{\bf g}(C(u))$.
  2. (2)
    $\mathit{mn}(v)\in{\bf h}_{C}(u)$.
Clearly, the childchild relation is completely determined by $\bf t$} and $({\bf C},{\bf D},{\bf K},{\bf f},{\bf g})$.
Lemma 3.
$(V(H),child)=F$.
Proof.
We first need to clarify that the connected components of each ${\bf g}(C)$ are, in fact stains of an arbitrary vertex of the component.
Claim 4.
For each $C\in{\bf C}$ and each $u\in C$, ${\bf h}_{C}(u)=\mathit{Stain}(u,({\bf t},F))$.
Proof.
As $u\in\mathit{Stain}(u,({\bf t},F))\subseteq V({\bf g}(C))$ and $\mathit{Stain}(u,({\bf t},F))$ is a connected subset of $T$, we conclude that $\mathit{Stain}(u,({\bf t},F))\subseteq{\bf h}_{C}(u)$. Assume that $\mathit{Stain}(u,({\bf t},F))\neq{\bf h}_{C}(u)$. Then there is $v_{0}\in\mathit{Stain}(u,({\bf t},F))$ and $v_{1}\notin\mathit{Stain}(u,({\bf t},F))$ such that ${\bf g}(C)$ has an edge between $v_{0}$ and $v_{1}$. By construction, the edge is an edge of $T[\mathit{Stain}(u^{\prime},({\bf t},F))]$ for some $u^{\prime}\in C$. We note that $u^{\prime}\neq u$ since $v_{1}\notin\mathit{Stain}(u,({\bf t},F))$. and that $v_{0}\in\mathit{Stain}(u,({\bf t},F))$. It follows the stains of $u$ and $u^{\prime}$ intersects but this is a contradiction to $C$ being an independent set, ∎
Assume now that $(u,v)\in F$. By definition, $v\in{\bf f}(C(u))$ and also $\mathit{mn}(v)\in\mathit{Stain}(u,({\bf t},F))$. By Claim 4, $v\in{\bf h}_{C}(u)$. We conclude that $(u,v)\in child$.
Conversely, assume that $(u,v)\in Child$. Then, by definition, $v$ is not a root of $T$. Hence, $parent(v),v)$ is an edge of $F$. We have already observed that $(parent(v),v)\in child$. We proceed to note that the in-degree of $v$ in child is exactly one, which will imply that $parent(v)=u$ and hence $(u,v)\in F$. Indeed, assume that there are two distinct $u_{1},u_{2}$ such that both $(u_{1},v)$ and $(u_{2},v)$ are elements of childchild. If $C(u_{1})\neq C(u_{2})$ then $v$ has two different parents in $F$, a contradiction, hence we conclude that $C(u_{1})=C(u_{2})=C$. Now, it follows from Claim 4 that $v\in\mathit{Stain}(u_{1},({\bf t},F))\cap\mathit{Stain}(u_{2},({\bf t},F))$ in contradiction to the independence of $u_{1}$ and $u_{2}$. ∎
The main idea of the proof of Lemma 9 is guessing a transduction tuple. A subtle aspect of the proof is guessing of $K$ that is a family of subgraphs of $T$ rather than a family of sets. The guessing is done by guessing two sets per subgraph. One of them is the set of vertices of the subgraphs and the other is a subset of the first set consisting of the vertices that are not connected to their parents in the resulting subgraph.
Proof of Lemma˜9.
The algorithm produces a transduction
${\bf Tr}=Tr[{\bf c}_{1},\dots,c_{q},{\bf c^{*}},{\bf d}_{0},\dots,{\bf d}_{q},{\bf d^{*}},{\bf k}^{0}_{1},{\bf k}^{1}_{1},\dots,{\bf k}^{0}_{q},{\bf k}^{1}_{q},{\bf k^{*}},{\bf x}_{0},{\bf x}_{1}]$
The elementary transductions ${\bf c}_{1},\dots,{\bf c}_{q}$ are all extensions introducing respective new unary relations $C_{1},\dots,C_{q}$. Let $\tau_{1}$ be the signature obtained from $\tau_{td}$ by adding $C_{1},\dots,C_{q}$. Then $\bf c^{*}$} is the domain filtering transduction whose input signature is $\tau_{1}$. The associated MSO sentence is true if and only if $C_{1},\dots,C_{q}$ weakly partition $Vertex$ (that is, one or more of $C_{i}$ may be empty sets).
The elementary transductions ${\bf d}_{0},\dots,{\bf d}_{q}$ are all extensions introducing respective new unary relations $D_{0},\dots,D_{q}$. Let $\tau_{2}$ be the signature obtained from $\tau_{1}$ by adding the relation $D_{0},\dots,D_{q}$. Then $\bf d^{*}$} is the domain filtering transduction whose input is $\tau_{2}$. The associated MSO sentence is true if and only $D_{0},\dots,D_{q}$ weakly partition $Vertex$.
The elementary transactions ${\bf k}^{0}_{1},{\bf k}^{1}_{1},\dots,{\bf k}^{0}_{q},{\bf k}^{1}_{q}$ are all extensions introducing respective new unary relations $K^{0}_{1},K^{1}_{1},\dots,K^{0}_{q},K^{1}_{q}$. Let $\tau_{3}$ be the signature obtained from $\tau_{2}$ by introducing the relations $K^{0}_{1},K^{1}_{1},\dots,K^{0}_{q},K^{1}_{q}$. Then $\bf k^{*}$} is the domain filtering transduction whose accompanying MSO sentence is true if and only the following statements hold.
  1. (1)
    Each $K_{i}^{j}$ is a subset of $Node$. Moreover, each $K_{i}^{1}$ is a subset of $K_{i}^{0}$. For each $1\leq i\leq q$, and each vertex $\in K^{1}_{i}, uu$ is not a root of $T$ and, moreover, $parent(u)$, the parent of $u$, belongs to $K^{0}_{i}$.
  2. (2)
    For each $1\leq i\leq q$, let us define $T_{i}$ as the subgraph of $T[K^{0}_{i}]$ obtained by removal from $T[K^{0}_{i}]$ the edges $(parent(u),u)$ for each $u\in K^{1}_{i}$. Then there is a bijection $\gamma_{i}$ from $C_{i}$ to the connected components of $T_{i}$ so that each $u\in C_{i}$ is mapped to the component containing $\mathit{mn}(u)$.
Next, ${\bf x}_{0}$ is an interpretation transduction mapping a model of $\tau_{3}$ to the model over $\tau_{ef}$ over the same universe.. Thus the underlying tuple of MSO formulas is $(iv(x),ie(x),ia(x),\varphi(x,y))$. The formulas $iv(x),ie(x),ia(x,y)$ are true if and only if, respectively $x\in Vertex$, $x\in Edge$ and $ia(x,y)$ is an element of $Adjacent$. In other words, these formulas simply copy $Vertex$, EdgeEdge, and $Adjacent$ relation to the new model.
The formula $\varphi(x,y)$ is is true if and only if the following statements hold.
  1. (1)
    Let $i$ be such that $y\in D_{i}$. Then $i>0$ and $x\in C_{i}$.
  2. (2)
    $\mathit{mn}(y)$ belongs to the connected component of $T_{i}$ that contains $\mathit{mn}(x)$.
Finally, ${\bf x}_{1}$ is a domain filtering transduction over signature $\tau_{ef}$ associated with a formula that is true over a model $M$ of $\tau_{ef}$ if and only if $F(M)$ is an elimination forest of $H(M)$.
We observe that the transduction as above is straightforward to be produced by an algorithm with input $q$, that the transduction size is linear in $q$ and that for each $M_{1}\in{\bf Tr}(M({\bf t},F))$, $M_{1}$ is well formed in the sense that the $Child$ relation of the model is a binary relation over $Vertex$.
It remains to demonstrate that, when $q\geq q_{0}$ where $q_{0}$ is the chromatic number of $\mathsf{CG}({\bf t},F)$ then ${\bf Tr}(M(H,{\bf t}))$ contains an output $M^{*}$ such that $F(M^{*})=F$
In this case, there is $M_{1}\in Tr[{\bf c}_{1},\dots,{\bf c}_{q}](M(H,{\bf t}))$ that is obtained by introduction of relations $C_{1},\dots,C_{q}$ such that $C_{1},\dots,C_{q_{0}}$ are colouring classes of $\mathsf{CG}({\bf t},F)$ and, if $q_{0}<q$, $C_{q_{0}+1},\dots,C_{q}$ are all empty sets. By construction, $M_{1}$ satisfies the underlying sentence of $\bf c^{*}$}. We conclude, therefore, that $M_{1}\in Tr[{\bf c}_{1},\dots\bf c_{q},{\bf c^{*}}](M(H,{\bf t}))$.
Next, let $D_{0}$ be the set consisting of the roots of $F$ and for each $1\leq i\leq q$, let $D_{i}$ consisting of all the children of the non-leaf elements of $C_{i}$. We observe that $D_{0},\dots D_{q}$ constitutes a weak partition of $V(H)$. Indeed, every non-root element appears in some $D_{i}$ due to $C_{1},\dots,C_{q}$ being a partition of $V(H)$. On the other hand, an element appearing in two distinct $D_{i}$ and $D_{j}$ means that either an element is both a root of $F$ and not a root of $F$ or that an element has two parents. Therefore, there is $M_{2}\in Tr[{\bf d}_{1},\dots,{\bf d}_{q}](M_{1})$ obtained from $M_{1}$ by adding relations $D_{0},\dots,D_{q}$ and $M_{2}$ satisfies the underlying sentence of $\bf d^{*}$}. In other words, $M_{2}\in Tr[{\bf d}_{1},\dots,{\bf d}_{q},{\bf d^{*}}](M_{1})$.
For each $i\in[q]$, let $K^{0}_{i}=\bigcup_{u\in C_{i}}\mathit{Stain}(u,({\bf t},F))$ and let $T_{i}=\bigcup_{u\in C_{i}}T[\mathit{Stain}(u,({\bf t},F))]$. We observe that there is $K^{1}_{i}\subseteq K^{0}_{i}$ that $T_{i}$ is obtained from $T[K^{0}_{i}]$ by removal of edges $\{(parent(u),u)|u\in K^{1}_{i}\}$. By definition, there is $M_{3}\in Tr[{k}^{0}_{1},{\bf k}^{1}_{1},\dots,{\bf k}^{0}_{q},{\bf k}^{1}_{q}](M_{2})$ obtained from $M_{2}$ by introduction of relations $K^{0}_{1},K^{1}_{1},\dots,K^{0}_{q},K^{1}_{q}$. By definition, $M_{3}$ satisfies the sentence $\bf k^{*}$} and hence
$M_{3}\in Tr[{k}^{0}_{1},{\bf k}^{1}_{1},\dots,{\bf k}^{0}_{q},{\bf k}^{1}_{q},{\bf k^{*}}](M_{2}).$
Let ${\bf C}=\{C_{1},\dots,C_{q_{0}}\}$. Let ${\bf D}=\{D_{i}|0\leq i\leq q_{0}\}$. Let ${\bf K}=\{T_{i}|i\in[q]\}$. Let $\bf f$} be a function on $\bf C$} that maps each $C_{i}$ to $D_{i}$. Let $\bf g$} be a function on $\bf C$} that maps each $C_{i}$ to $T_{i}$.
We observe that $({\bf C},{\bf D},{\bf K},{\bf f},{\bf g})$ is a transduction tuple of $({\bf t},F)$. Furthermore, the relation determined by $\varphi(x,y)$ of ${\bf x}_{0}$ is the childchild relation for $({\bf C},{\bf D},{\bf K},{\bf f},{\bf g})$ as in Definition 2. It follows from Lemma 3 that there is $M^{*}\in Tr[{\bf x}_{0}](M_{3})$ such that $F(M^{*})=F$. As $F$ is a proper elimination forest of $H$, $M^{*}\in Tr[{\bf x}_{0},{\bf x}_{1}](M_{3})$ thus completing the proof of the lemma. ∎
Appendix E Proofs Details for Section 4
In this section we provide full proofs of the statements in Section 4.
Proof of Corollary 5.
Just run the algorithm for the testing whether the $\mathcal{F}_{\delta}$-width is at most $k$ as stated in Theorem 2. Suppose that the algorithm rejects. As $\mathcal{F}_{\delta}\subseteq\mathcal{F}^{*}_{\delta}$, also $\mathcal{F}_{\delta}$-$\leq\mathcal{F}^{*}_{\delta}-width(H)(H)$, it is guaranteed that the latter is greater than $k$. Assume now that the algorithm accepts. For the sake of contradiction, let us suppose that the $\mathcal{F}^{*}_{\delta}$-width of $H$ is greater than $2k$. This means that there is a specific $f\in\mathcal{F}^{*}_{\delta}$ such that the $f$-width of $H$ is greater than $2k$. Let $f_{0}$ be a function obtained from $f$ as follows: for each $v\in V(H)$ if $f(v)=a/\delta$ then $f_{0}(v)=\frac{\lfloor a\rfloor}{\delta}$. Recall that by definition of $\mathcal{F}^{*}_{\delta}$, $a\geq 1$ and therefore $a\leq 2\lfloor a\rfloor$. Since the initial algorithm accepted, the $f_{0}$-width of $H$ must be at most $k$. Let $(T,B)$ be a minimal $f_{0}$-width tree decomposition of $H$. Since for each $v\in V(H)$, $f(v)\leq 2f_{0}(v)$, we conclude that for each $u\in V(T)$, $f(B(u))\leq 2f_{0}(B(u))\leq 2k$ meaning that $f$-width of $H$ is at most $2k$ in contradiction to our assumption. ∎
Proof of Theorem 2.
With the notation as in Theorem 3, let $\psi=\forall U_{0}\dots\forall U_{\delta}\varphi(U_{0},\dots,U_{\delta})$.
Claim 1.
$M\models\psi$ if and only if $\mathcal{F}_{\delta}$-width of $H$ is at most $k$.
Proof of Claim. Assume first that $M\models\psi$. For the sake of contradiction, assume that the $\mathcal{F}_{\delta}$-width of $H$ is greater than $k$. This means that there is a function $f\in\mathcal{F}_{\delta}$ such that the $f\textsf{-width}$ of $H$ is greater than $k$. By definition of $\mathcal{F}_{\delta}$, there is a weak partition $U^{\prime}_{0}\dots U^{\prime}_{\delta}$ such that for each $i\in\{0,\dots\delta\}$ $U^{\prime}_{i}$ consists of exactly those elements $v$ such that $f(v)=\frac{i}{\delta}$. Clearly, $f=f[U^{\prime}_{0},\dots U^{\prime}_{\delta}]$. By definition $M\not\models\varphi(U^{\prime}_{0},\dots U^{\prime}_{\delta})$. Then $M\not\models\psi$ as witnessed by $U_{0}=U^{\prime}_{0},\dots,U_{\delta}=U^{\prime}_{\delta}$ thus contradicting our assumption.
Assume now that $M\not\models\psi$. This means that there are $U_{0}=U^{\prime}_{0},\dots U_{\delta}=U^{\prime}_{\delta}$ such that $M\not\models\varphi(U^{\prime}_{0}\dots,U^{\prime}_{\delta})$. By definition, this means that $f=f[U^{\prime}_{0},\dots,U^{\prime}_{\delta}]$ is an independent set function of $H$ and the $f\textsf{-width}$ of $H$ is greater than $k$. By definition of $f$, clearly $f\in\mathcal{F}_{\delta}$. It follows that $\mathcal{F}_{\delta}$-width of $H$ is greater than $k$. ∎
In order to design the algorithm for testing $\mathcal{F}_{\delta}$-width, we observe first that $tw(H)$ is at most twice the $\mathcal{F}_{\delta}$-width multiplied by $rank(H)$. In fact, we can prove this using the $\mathcal{F}_{\delta}^{*}$-width instead of $\mathcal{F}_{\delta}$-width. Indeed, $f$ be the function assigning each vertex with weight $1/rank(H)$ (note that this requires $\delta\geq rank(H)$). Clearly, $f$ is an independent set function. Furthermore, it is not hard to observe that $tw(H)$ is at most the $f\textsf{-width}$ of $H$ multiplied by $rank(H)$ whereas the $f\textsf{-width}$ of $H$ is at most the adaptive width of $H$. Then the bound follows using Corollary˜5.
The first step of the algorithm is running a linear time algorithm for treewidth computation parameterized by $k\cdot\delta$. If the treewidth construction algorithm rejects then our algorithm can safely return $False$ by the previous paragraph. Otherwise, the treewidth construction algorithm returns a tree decomposition ${\bf t}=(T,{\bf B})$ of treewidth at most $k\cdot\delta$. We run $\mathcal{A}(H,{\bf t},k,\delta)$ as in the statement of Theorem 3. We note that, by definition of $\bf t$}, the runtime of the algorithm $\mathcal{A}$ is FPT parameterized by $k,rank(H),\Delta(H)$, and $\delta$. Note that it follows that the size of $\varphi(U_{0},\dots,U_{\delta})$ is also upper bounded by a function depending on these parameters. Clearly, that the runtime of construction of $\psi$ is FPT parameterized by $k,rank(H),\Delta(H)$, and $\delta$ and the size of $\psi$ is upper-bounded by a function depending on these parameters.
By Claim 1, it only remains to test whether $(H,{\bf t})\models\psi$. This can be done within the required runtime according to Courcelle’s theorem. ∎
In the rest of this section we prove Theorem 3. Let us start from an easy step: exclusion of the case where $U_{0},\dots,U_{q}$ do not form a weak partition of $V(H)$. This can be done by conjunction of two formulas: one being true when the weak partition does not hold and the other as specified in the theorem but assuming the weak partition to hold. The first formula is easy to implement so in what follows we will prove the theorem under assumption that $U_{0},\dots U_{\delta}$ do form a partition of $V(H)$.
We next note that the function $f[U_{0},\dots,U_{\delta}]$ does not have the bounded size property. Indeed, as vertices of $U_{0}$ do not increase the value of the function, a function can have small or even zero value on a set that is a superset of $U_{0}$. We overcome this obstacle by proving Theorem 3 for functions where $U_{0}$ is assumed to be empty. The MSO resulting from the algorithm as specified in the restricted version in then‘wrapped’ into a transformation that applies it to the submodel of the input model where the universe is restricted to exclude the values of $U_{0}$.
Towards the implementation of this plan let the sets $U_{1},\dots,U_{\delta}$ be a partition of $V(H)$ and let $f_{1}=f_{1}[U_{1},\dots,U_{\delta}]$ be a function that assigns each vertex of $U_{i}$ with weight $\frac{i}{\delta}$, As we specified earlier, $f_{1}$ can also be seen as a width function mapping a subset to the sum of weights of its elements. Now, we reformulate Theorem 3 for functions $f_{1}[U_{1},\dots,U_{\delta}]$
Theorem 2.
There is an algorithm $\mathcal{A}(H,{\bf t},k,\delta)$ (where $H$ is a hypergraph, $\bf t$} is a tree decomposition of $H$, $k$ is a non-negative rational, and $\delta\in\mathbb{N}$) that returns an MSO formula $\varphi_{1}(U_{1},\dots,U_{\delta})$ where $U_{0},\dots,U_{\delta}$ are free set variables so that the following holds under assumption that $U_{1},\dots,U_{\delta}$ weakly partition $H$.
Let $M=M(H,{\bf t})$ be a model over $\tau_{td}$. Then the following statements hold.
  1. (1)
    If $f_{1}[U_{1},\dots,U_{\delta}]$ is not an independent set function then $M\models\varphi_{1}(U_{1},\dots,U_{\delta})$
  2. (2)
    If $f_{1}=f_{1}[U_{1},\dots,U_{\delta}]$ is an independent set function and and the $f_{1}$-width of $H$ is at most $k$ then $M\models\varphi_{1}(U_{1},\dots,U_{\delta})$.
  3. (3)
    If $f_{1}=f_{1}[U_{1},\dots,U_{\delta}]$ is an independent set function and and the $f_{1}$-width of $H$ is greater than $k$ then $M\not\models\varphi_{1}(U_{1},\dots,U_{\delta})$.
The runtime of the algorithm is FPT parameterized by the treewidth of $\bf t$}, $k$, $rank(H)$, $\Delta(H)$ and $\delta$ and the size of the resulting formula is upper-bounded by a function depending on these parameters.
Now, we carry out the MSO transformation. Let $M$ be a model over some signature $\tau$ and let $U_{0}$ be a subset of the universe $U(M)$ of $M$. The model $M\setminus U_{0}$ is also over $\tau$, has the universe $U(M)\setminus U_{0}$ and each relation $R$ of $M$ is replaced with $R\setminus U_{0}$ obtained from $R$ by removal of all tuples containing an element of $U_{0}$.
Lemma 3.
Let $\psi_{0}(V_{1},\dots,V_{r})$ be an MSO formula with $V_{1},\dots,V_{r}$ being free set variables. Then there is an algorithm that transforms $\psi_{0}(V_{1},\dots,V_{r})$ to an MSO formula $\psi_{1}(V_{0},\dots,V_{r})$ over the same signature $\tau$ where $V_{0}$ is also a free set variable so that the following holds. Let $V^{\prime}_{0},\dots,V^{\prime}_{r}$ be a respective instantiation to $V_{0},\dots,V_{r}$. Then for each model $M$ over $\tau$, $M\models\psi_{1}(V^{\prime}_{0},\dots,V^{\prime}_{r})$ if and only if $M\setminus V^{\prime}_{0}\models\psi_{0}(V^{\prime}_{1},\dots,V^{\prime}_{r})$.
Proof.
In the formula $\psi_{0}(V_{1},\dots,V_{r})$ accompany each quantifier over the individual variables with a ‘guard’ to ensure that only the variables outside of $V_{0}$ are considered. In particular $\exists vX$ is replaced with $\exists v(v\notin V_{0})\wedge X$ and $\forall vX$ is replaced with $\forall v(v\notin V_{0})\rightarrow X$. The correctness can be verified by a straightforward bottom up inductive verification on the structure of the formula. ∎
Now, we are ready to prove Theorem 3.
Proof of Theorem 3.
Run the algorithm $\mathcal{A}(H,{\bf t},k,\delta)$ as in Theorem 2 to return the MSO formula $\varphi_{1}(U_{1},\dots,U_{\delta})$ as stated in Theorem 2. Let $\varphi(U_{0},\dots,U_{\delta})$ be obtained from $\varphi_{1}(U_{1},\dots,U_{\delta})$ by the transformation as in Lemma 3.
Let $U^{\prime}_{0},\dots,U^{\prime}_{\delta}$ be a respective instantiation to $U_{0},\dots,U_{\delta}$ that constitutes a weak partition of $V(H)$. Suppose that $M\models\varphi(U^{\prime}_{0},\dots,U^{\prime}_{\delta})$. By Lemma 3, $M\setminus U^{\prime}_{0}\models\varphi_{1}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$. There may be two reasons for that. One reason is that $f_{1}=f_{1}[U^{\prime}_{1},\dots,U^{\prime}_{\delta}]$ is not an independent set function of $H\setminus U^{\prime}_{0}$. This means that $H$ has a hyperedge $e$ such that $f_{1}(e\setminus U^{\prime}_{0})>1$. Clearly, this means that $f(e)>1$ where $f=f[U^{\prime}_{0},\dots,U^{\prime}_{\delta}]$. We conclude that $f$ is not an independent set function. The second reason that $M\setminus U^{\prime}_{0}\models\varphi_{1}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ is that $f_{1}$ is an independent set function and the $f_{1}$-width of $H\setminus U^{\prime}_{0}$ is at most $k$. Let $(T,{\bf B})$ be a tree decomposition of $H\setminus U^{\prime}_{0}$ of $f_{1}$-width at most $k$. Let $\bf B^{*}$} be obtained from $\bf B$} by setting ${\bf B^{*}}(t)={\bf B}(t)\cup U^{\prime}_{0}$ for each $t\in V(T)$. It is not hard to see that $(T,{\bf B^{*}})$ is a tree decomposition of $H$ of $f$-width at most $k$.
Assume now that $M\not\models\varphi(U^{\prime}_{0},\dots,U^{\prime}_{\delta})$. By Lemma 3, $M\setminus U^{\prime}_{0}\not\models\varphi_{1}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$. The only reason for that is $f_{1}$ is an independent set function and that $f_{1}$-width of $H\setminus U^{\prime}_{0}$ is greater than $k$. It follows that $f$ is an independent set function of $H$. For the sake of contradiction, assume that the $f$-width of $H$ is at most $k$. Let $(T,{\bf B^{*}})$ be a tree decomposition of $H$ of $f$-width at most $k$. Let $\bf B$} be obtained from $\bf B^{*}$} by setting ${\bf B}(t)={\bf B^{*}}\setminus U^{\prime}_{0}$ for each $t\in V(T)$. It is not hard to see that $(T,{\bf B})$ is a tree decomposition of $H\setminus U^{\prime}_{0}$ of $f_{1}$-width at most $k$ in contradiction to our assumption. We thus conclude that indeed $f$-width of $H$ is greater than $k$. It follows from the previous discussion that $\varphi(U_{0},\dots,U_{q})$ is as specified in the statement of the theorem. ∎
It thus remains to prove Theorem 2.
Proof of Theorem 2.
Let $U_{1},\dots,U_{q}$ be a weak partition of $V(H)$ Let $U\subseteq V(H)$. We refer to $(|U\cap U_{1}|,\dots,|U\cap U_{\delta}|)$ as the partition tuple of $U$ (w.r.t. $U_{1},\dots,U_{\delta}$ if not clear from the context). The following easy to establish claim is stated without a proof.
Claim 4.
For tuple $p_{1},\dots,p_{\delta}$ of non-negative integers, there is an MSO formula $\varphi_{p_{1},\dots,p_{\delta}}(U,U_{1},\dots,U_{\delta})$ where $U,U_{1},\dots,U_{\delta}$ are free set variables that is satisfied for the respective instantiation $U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ if and only if $(p_{1},\dots,p_{\delta})$ is the partition tuple of $U^{\prime}$ w.r.t. $U^{\prime}_{1},\dots,U^{\prime}_{\delta}$. The size of the formula is upper-bounded by a function of $p_{1}+\dots+p_{\delta}$.
This leads us stating existence of an MSO formula that will be an important building block for the algorithm being designed.
Claim 5.
For a positive rational $p$, there is an MSO formula $\varphi_{\delta,p}(U,U_{1},\dots,U_{\delta})$ where $U,U_{1},\dots,U_{\delta}$ are free set variables so that the following holds. Let $U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ be a respective instantiation for $U,U_{1},\dots,U_{\delta}$. Then $\varphi_{\delta,p}(U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ is satisfied if and only $f[U^{\prime}_{1},\dots,U^{\prime}_{\delta}](U^{\prime})$ is at most $p$. The size of the formula is upper bounded by a function depending of $p$ and $\delta$.
Proof of Claim. Let ${\bf T}_{\delta,p}$ be the set of all tuples $(p_{1},\dots,p_{\delta})$ of non-negative integers such that $\sum_{i\in[\delta]}\frac{i\cdot p_{i}}{\delta}\leq p$. Clearly the number of such tuples is upper bounded by a function depending on $p$ and $\delta$. We set $\varphi_{\delta,p}(U,U_{1},\dots,U_{\delta})=\bigvee_{(p_{1},\dots,p_{\delta})\in{\bf T}_{\delta,p}}\varphi_{p_{1},\dots,p_{\delta}}(U,U_{1},\dots,U_{\delta})$, where $\varphi_{p_{1},\dots,p_{\delta}}(U,U_{1},\dots,U_{\delta})$ is as in Claim 4.
Let $U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ be an a respective instantiation to the free variables. Assume that $\varphi_{\delta,p}(U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ is satisfied. Then there is $(p_{1},\dots,p_{\delta})\in{\bf T}_{\delta,p}$ such that $\varphi_{p_{1},\dots,p_{\delta}}(U,U_{1},\dots,U_{\delta})$ is satisfied. By Claim 4, the partition tuple of $U^{\prime}$ w.r.t. $U^{\prime}_{1},\dots U^{\prime}_{\delta}$ is precisely $(p_{1},\dots,p_{\delta})$. We note that $f[U^{\prime}_{1},\dots,U^{\prime}_{\delta}](U^{\prime})$ is precisely $\sum_{i\in[\delta]}\frac{i\cdot p_{i}}{\delta}$. As $(p_{1},\dots,p_{\delta})\in{\bf T}_{\delta,p}$, this quantity is at most $p$.
Assume now that $\varphi_{\delta,p}(U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ is not satisfied. For the sake of contradiction, assume that $f[U^{\prime}_{1},\dots U^{\prime}_{\delta}](U^{\prime})\leq p$. As we observed in the previous paragraph, this means that the partition tuple $(p_{1},\dots,p_{\delta})$ of $U^{\prime}$ w.r.t. $U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ is an element of ${\bf T}_{\delta,p}$. Consequently, $\varphi_{p_{1},\dots,p_{\delta}}(U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ is one of the disjuncts and it is satisfied according to Claim 4 in contradiction to non-satisfaction of $\varphi_{\delta,p}(U^{\prime},U^{\prime}_{1},\dots,U^{\prime}_{\delta})$.
It remains to add that $p_{1}+\dots+p_{\delta}$ is at most $p\delta$ and hence the size upper bound is satisfied according to Claim 4. ∎
We use $\varphi_{\delta,p}(U,U_{1},\dots,U_{\delta})$ for testing that $f_{1}=f_{1}[U_{1},\dots U_{\delta}]$ is an independent set function and for testing whether the $f_{1}$-width of $H$ is at most $k$. The exact statement is provided in the next claim.
Claim 6.
The following two statements hold:
  1. (1)
    There is an MSO formula $IsIndepSet(U_{1},\dots,U_{\delta})$ over $\tau_{td}$ so that the following holds. Let $U^{\prime}_{1},\dots U^{\prime}_{\delta}$ be a respective instantiation to $U_{1},\dots,U_{\delta}$ and let $M$ be a model for $\tau_{td}$.
    Then $M\models IsIndepSet(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ if and only if $f[U^{\prime}_{1},\dots U^{\prime}_{\delta}]$ is an independent set function for $H$.
  2. (2)
    There is a function $IsWidth_{k}(U_{1},\dots U_{\delta})$ over $\tau_{ef}$ so that the following holds. Let $U^{\prime}_{1},\dots U^{\prime}_{\delta}$ be a respective instantiation to $U_{1},\dots,U_{\delta}$ and let $M$ be a model over $\tau_{ef}$.
    Then $M\models IsWidth_{k}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ if and only if $f_{1}$-width of $F(M)$ is at most $k$.
The size of both formulas is upper bounded by a function depending on $k$ and $\delta$.
Proof of Claim. We define $IsIndepSet(U_{1},\dots,U_{\delta})$ as $\forall U~IsEdge(U)\rightarrow\varphi_{\delta,1}(U,U_{1},\dots,U_{\delta})$, where $IsEdge(U)$ is true if and only if $U\in E(H)$. The design of $IsEdge(U)$ is elementary and we do not provide further details. It follows from Claim 5 that for an instantiation $U^{\prime}_{1},\dots U^{\prime}_{\delta}$, $M\models IsIndepSet(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ if and only if $f_{1}[U^{\prime}_{1},\dots U^{\prime}_{\delta}]$ is at most $1$ for each hyperedge of $H$, exactly as in the first statement of the lemma.
We define $IsWidth_{k}(U_{1},\dots,U_{\delta})$ as $\forall U~IsBag(U)\rightarrow\varphi_{\delta,k}(U,U_{1},\dots,U_{\delta})$, where $IsBag(U)$ is as defined in the proof of Lemma 11. It follows from Claim 5 that for an instantiation $U^{\prime}_{1},\dots U^{\prime}_{\delta}$, $M\models IsWidth_{k}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$ if and only if $f_{1}[U^{\prime}_{1},\dots U^{\prime}_{\delta}]$ is at most $k$ for each bag of the tree decomposition induced by $F(M)$ thus confirming the second statement of the claim.
It remains to notice that it follows from Claim 5 that the size of both formulas is upper-bounded by a function dependent on $k$ and $\delta$. ∎
We define $\varphi_{1}(U_{1},\dots U_{\delta})$ as $IsIndepSet(U_{1},\dots,U_{\delta})\rightarrow\varphi_{2}(U_{1},\dots,U_{\delta})$ where $\varphi_{2}(U_{1},\dots,U_{\delta})$ is like $\varphi_{1}(U_{1},\dots,U_{\delta})$ but under assumption that $f_{1}[U_{1},\dots,U_{\delta}]$ is an independent set function. The design of $\varphi_{2}(U_{1},\dots,U_{\delta})$ essentially uses the reasoning for the manageable functions but with two nuances. The first is that the last element of the transduction is now $IsWidth_{k}(U_{1},\dots,U_{\delta})$ and the second nuance is that the transduction is transformed to an MSO formula rather than to an algorithm. We are now going to address the second nuance.
First of all, in light of the proof of Theorem 3 and, in particular, Lemma 9, we only need to consider transductions involving elementary transductions of colouring, filtering, and interpretation. Second, the last elementary transduction in the transduction being constructed is a filtering transduction associated with the formula $IsWidth_{k}(U_{1},\dots U_{\delta})$. We thus define a transduction $Tr[{\bf q}_{1},\dots,{\bf q}_{m}](Model,U_{1},\dots U_{\delta})$ as follows. Let $M,U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ be a respective instantiation to $Model,U_{1},\dots U_{\delta}$. Let ${\bf M}=Tr[{\bf q}_{1},\dots,{\bf q}_{m-1}](M)$. Then the output of $Tr[{\bf q}_{1},\dots,{\bf q}_{m}](M,U^{\prime}_{1},\dots U^{\prime}_{\delta})$ is $\{M^{\prime}|M^{\prime}\in{\bf M},M^{\prime}\models IsWidth_{k}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})\}$.
Claim 7.
There is an algorithm that transforms $Tr[{\bf q}_{1},\dots,{\bf q}_{m}](Model,U_{1},\dots U_{\delta})$ into a MSO formula $\varphi^{*}(U_{1},\dots,U_{\delta})$ over the signature as that of $Model$ so that the following holds. Let $M,U^{\prime}_{1},\dots,U^{\prime}_{\delta}$ be a respective instantiation to $Model,U_{1},\dots U_{\delta}$. Then $M\models\varphi^{*}(U^{\prime}_{1},\dots,U^{\prime}_{\delta})$
if and only if $Tr[{\bf q}_{1},\dots,{\bf q}_{m}](M,U^{\prime}_{1},\dots U^{\prime}_{\delta})\neq\emptyset$.
Proof of Claim. For the purpose of being used as data in an algorithm, a transduction can be represented as a structure of size proportional to the size of the transduction (as per Proposition 12). An elementary transduction can be represented as $(type,sequence)$ where typetype is the transduction type and sequencesequence is a possibly empty sequence of accompanying MSO formulas. The colouring transduction is presented as $(colour,())$, The filtering transduction is represented as $(filter,(\varphi^{\prime}))$ where $\varphi^{\prime}$ is the MSO formula accompanying the filtering transduction. Finally, the interpretation is represented as $(interpret,{\bf Rel})$ where $\bf Rel$} is the sequence of $\varphi_{R}$ for each relation in the input signature of the transduction. The transduction $Tr[{\bf q}_{1},\dots\bf q_{a}]$ is represented as the transduction sequence $(rep({\bf q}_{1}),\dots,rep({\bf q}_{a}))$ where $rep({\bf q}_{i})$ is the representation of ${\bf q}_{i}$. We demonstrate turning a transduction into an MSO formula by induction on the length of the transduction sequence.
If the length is $1$ then the only elementary transduction of the sequence is the filtering transduction accompanied with the formula $IsWidth_{k}(U_{1},\dots U_{\delta})$. Then this formula is the output of the transformation.
Assume now that the transduction sequence is $(rep({\bf q}_{1}),\dots,rep({\bf q}_{a}))$ with $a\geq 2$. By the induction assumption, $(rep({\bf q}_{2}),\dots,rep({\bf q}_{a}))$ can be transformed into an MSO formula $\varphi^{*}(U_{1},\dots,U_{\delta})$ for the same signature as that of $rep({\bf q}_{2})$ with the property as specified in the statement of the claim.
Assume first that ${\bf q}_{1}$ is a colouring transduction. This means that the signature of ${\bf q}_{2}$ is obtained from the signature of ${\bf q}_{1}$ by adding one unary relation $U$. We note that $\varphi^{*}(U_{1},\dots,U_{\delta})$ can be thought as the formula $\varphi^{*}(U,U_{1},\dots,U_{\delta})$ for the signature of ${\bf q}_{1}$ where $U$ is considered a free variable. We define the MSO representation for $(rep({\bf q}_{1}),\dots,rep({\bf q}_{a}))$ as $\exists U\varphi^{*}(U,U_{1},\dots,U_{\delta})$.
Assume next that ${\bf q}_{1}$ is a filtering transduction accompanied with an MSO formula $\varphi^{\prime}$. We define the MSO representation for $(rep({\bf q}_{1}),\dots,rep({\bf q}_{a}))$ as $\varphi^{\prime}\wedge\varphi^{*}(U_{1},\dots,U_{\delta})$.
Finally, we assume that ${\bf q}_{1}$ is the interpretation transduction. This means that each relation $R$ of the signature of ${\bf q}_{2}$ corresponds to the formula $\varphi_{R}$ of the transduction having the same arity as $R$. We define the MSO representation for $(rep({\bf q}_{1}),\dots,rep({\bf q}_{a}))$ by replacing each occurrence of $R(x_{1},\dots,x_{b})$ in $\varphi^{*}(U_{1},\dots,U_{\delta})$ by $\varphi_{R}(x_{1},\dots,x_{b})$.
The correctness of the transformation can be verified by a straightforward but tedious verification by induction on the length of the transduction sequence. We omit the details. ∎
Claim 8.
There is a monotone function $\alpha^{*}=\alpha^{*}(tw({\bf t}),k,rank(H),\Delta(H),\delta)$ such that if $f_{1}$-width of $H$ is at most $k$ then there is an elimination forest of $f_{1}$-width of $H$ such that the chromatic number of $\mathcal{CG}({\bf t},F)$ is at most $\alpha^{*}$.
Proof of Claim. We note that the bag size for a tree decomposition of $H$ of $f_{1}$-width at most $k$ is at most $k\delta$. Also there are at most $k\delta+1$ distinct values of $f_{1}$-width that are at most $k$. By Lemma 7 this means that, under the premises of the claim, there is an elimination forest $F$ of $H$ such that both splitsplit and mirmir measures of $({\bf t},F)$ are upper bounded by a monotone function depending on the parameters as those of $\alpha^{*}$. The claim is now immediate from Theorem 10. ∎
We now have all the pieces to design the algorithm as in the statement of Theorem 3. The algorithm consists of the following steps.
  1. (1)
    Compute an upper bound $\alpha^{*}$ as in Claim 8
  2. (2)
    Compute the transduction $\bf Tr$} of length $\alpha^{*}$ as in Theorem 9.
  3. (3)
    Append $IsWidth_{k}(U_{1},\dots,U_{\delta})$ to the end of $\bf Tr$} and let $\bf Tr^{*}$} be the resulting transduction.
  4. (4)
    Turn $\bf Tr^{*}$} into an MSO formula $\varphi_{2}(U_{1},\dots,U_{\delta})$ as per Claim 7
Let us prove correctness of the algorithm. Assume first that $f_{1}$-width of $H$ is at most $k$. By combination of Claim 8 and Theorem 9 an elimination forest $F$ of $f_{1}$-width at most $k$ belongs to the output of $\bf Tr$}. It follows from Claim 6 that the output of $\bf Tr^{*}$} is non-empty and it follows from Claim 7 that $M(H,{\bf t})\models\varphi_{2}(U_{1},\dots,U_{\delta})$.
Assume now that the $f_{1}$-width of $H$ is greater than $k$. As for each output $M^{\prime}$ of ${\bf Tr}(M(H,{\bf t}))$ $F(M^{\prime})$ is an elimination forest of $H$, the $f_{1}$-width of each $F(M^{\prime})$ is greater than $k$. It follows from Claim 6 that the output of ${\bf Tr^{*}}(M(H,{\bf t}))$ is empty. It then follows from Claim 7 that $M(H,{\bf t})\not\models\varphi_{2}(U_{1},\dots,U_{\delta})$. ∎